A Complete Bibliography of Publications in Communications in Statistics: Theory and

Methods: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.02

Title word cross-reference

16 [Dow76]. 18 [Dow76]. 2 [Sea79, Yam77]. 2^{m} [SG77]. 2 × 2 [KT78, Ray76]. 2 × j × k [Wyn77]. 3 [Yam77]. 4 ≤ m ≤ 8 [SG77]. A [Hen76]. C(α) [Ray76]. χ^{2} [Boa77, DS76b, McC78, HR76]. δ^{∗} [San76]. Δ_{P} [PS77a]. F [Amo76, MCCC76, Gor78]. F (x) [UW79]. g [Gor78]. G(y) = F (x − c) [UW79]. K [KR77, DNV77, Shi78]. M [Hil79]. Pr(y < x) [UW79]. t [RF79, Res76, SS79a]. U [Kre78, MS78a, Yam77]. w [McL77]. x ≥ y [HK76b].

-estimates [Hil79]. -goodness-of-fit [HR76]. -optimal [San76]. -optimality [Che78]. -statistics [Kre78, MS78a, Yam77]. -superior [PS77a]. -way [Sea79].

absolute [GKS77, Har77]. accelerated [Man78]. acceleration [PZW79]. acceptance [Spu78]. achieved [ABD77]. acreages [OB76]. Actual [DGL79]. Adaptive [HMC79, DV79, Haa78, Jon77a]. Adaptively [Ski78]. additive [Mie79c]. adequacy [LP77, Ste79]. Adjusting [SW77a]. affects [CLB79]. against [Adi76b, BY78]. age [Dow76]. Akaike [Sug78]. algebras [GM78]. algorithm [STT78]. algorithms [Scl77]. allocation [HR79]. allow [Spu78, SG77]. Alternative [MR79, Sea79, CI76, DH78, Gor78, RN79]. alternatives [Adi76b, BY78, Mar78a, SW77b]. amendments [Tam77a]. among [KR77]. Analysis [DM79, Gbu79, HH79, ISF79, PF77, RN79, SB76a, BT79, BSL79, CI76, Di 76, Di 77, Dip79, DG78, FH76, Gan78, Gar78, GH77, Gup79b, HRT77, HM79, Hui79, KG78, MH78, PR78, SK77, Sen79, SHW77, SH78, Zer79]. analysts [Sug78]. Analyzing [MSVS79, GK78]. Annex [Nak79]. ANOVA [Tam77a, Tam77b]. Application [SP76a, BT79, Di 76, DS76b, Gar78, GW77, Koo78, Ram77b, Smi79]. Applications [Ton77, Bre78, Di 77, GS78, Gup77a, MS78a, Ode76]. approach [Ban78a, DG79, GK78, Haa78, Kie78, Led79, Lin79, Lyn78, PS79, RN79, Rea77, Ste79, WP77]. approaches [BSL79]. approximate [Che79, Sch76]. approximation [HN78, Hid78, Ima76, MK76]. Approximations [MCCC76, ID76, Rob77]. areas [TG79a]. arising [AG78]. ascent [MK79]. aspects [TG79b]. assay [DM79]. assays [SD76]. assignment [Wei78a]. associated [BM78, TB79]. association [Agr77, AL79, PF77, Wil76]. assumption [SRC78]. asymmetric [AT79]. Asymptotic [GS78, Kie78, Lep76, Adi76a, AL79, AGS77, CK76, Ked77, KEKL76, LS79, Mie79a, Sch77]. asymptotically [Per77]. augmentation [CSH77]. autoregressive [Pee79]. auxiliary [DS77b]. average [Aro76, McC78, Pee79, SM76, SM77]. Ax [BK78b].

Balanced [PN76, PN77a, PN78, SG77, AF76, LE77, PF77, PN77b, SH78]. bank [Adi76a]. based [Agr77, BK79, Don77, FJ77, HL76, KA79, Koz79, LO76, LS79, MR79, MH76, Mic76, SS76, SS79b, Smi77a, Swi76, YM76]. basic [EHB77]. Basu [TB76]. Bayes [Ban78c, Ber79, HK76b, O’B79, PW77, PS79, SO79, SP76b, Yam77]. Bayesian [Day77, DG78, DG79, HB77, Leo77, Lin79, Mie79c, Oma78, SG76, Ste79, WB79]. behavior [DKH79]. Behrens [SRC78]. Bernoulli [Pri76]. best [DNV77, KW77, Ram77a, Rao78a]. beta [Hen76]. better [BM77, Kou77, Kou78]. between [GT76a, HRT77, NB78, Oli78, Wil76, Yam77, ZS76]. Bhattacharyya [DM76]. bias [AGS77, Di 76, Di 77, KC77, KC79, RW77, Seh78]. Biased [Dip79, Gup79c]. biases [BM78]. biasing [HK76a]. bibliography [HBB^{+}79]. bimodal [Cho78]. binary [PF77]. binomial [BM77, GT76a, Kor77, Nor76, Sch76, ZS76]. biological [Gan78]. biology [Koo78]. Bivariate [Nai76, Ave79, BS78, Bru79, Bru78, Dah77, Koz79, Kri76, Mui76, Pet79c, Smi77a, WLCC76]. block [HN78, HA78, PN77a, SP76a, SW77b, Ski78]. blocks [Zer79]. blues [Puk77]. Board [Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano76p, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano78o, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m]. Bonferroni [AS77]. bound [Koo78, TB79]. Bounds [AG78, BH78, SL79, Man78, Mar78a, PW78, UW79]. branching [Ram77b]. brief [Gov76, Gup77b]. Broemeling [Ton79].

calculating [DS76b]. Calculation [Man78]. calculations [Nak79]. canonical [Jam79]. Carlo [DL77, WC77, WLCC76]. case [APR79, ABC79, GD76, LE77, Pad77, SH78, ZS76]. categorical [Cli76, GK78, ISF79]. CD [ABD77]. cdf [Mic76, Smi77a]. censored [BT79, Dav78, Kam78, Lin79, MJB77, MJB79, PS79, SS79b, SB76b, TS77]. censoring [MS77, MS78b]. censorship [KR77]. census [CM77, KEKL76]. central [Amo76]. certain [DGL79, Ram77b, RY77, Wyn77]. chain [BiK79, FM78, HA78]. chance [KT78]. changing [Bro77a]. Characterization [BK78a, Ber78, Cho78, Fie77a, Gup79a, SS76, Wyn77]. Characterizations [Pee79, PR77]. characterizing [Kor77]. Chernoff [Hwa78]. chi [Hut79, Kri76, NS76, Rea77]. chi-square [NS76, Rea77]. chi-squared [Hut79, Kri76]. Choice [Cam79b, Rao78a]. Choosing [Rol76, BM77]. circular [Smi77a]. class [Ber78, GS78, HN79, HU77, Koc78, Lep77, Pol79, Rus76, Sha77, Wei78a]. classical [Mar78b, PR77, RW77]. classification [GG77, McL77, Sea79]. classifications [MBJ76]. classroom [Gen76]. clinical [Nor76, SS79b]. cloud [MSVS79, Smi79]. cloud-seeding [MSVS79, Smi79]. clustering [Scl77]. coat [HA78]. coat-of-mail [HA78]. Cochran [BN76]. coefficient [Agr77, HN79, Kha78, KR79, Sen78a, Wil78]. coefficients [Far78, HB78, SRC78, SM76, SM77]. column [Rus76]. combinations [DM76, DS76a, GW78, YM76]. combinatorial [Wyn77]. combining [SW77b, Ski78]. Comment [Haw77, Tik77, Far78]. comments [Cha76, Muk79]. common [MG79, Shi78]. Communications [GH78]. compared [DL77]. comparing [Tur76]. comparison [AS77, DK79, FPS77, HL76, KF78, Khu78, McL79, Oma78]. Comparisons [CK76, Tam77a, AM79, Bea77, BS76, EHB77, KR77, Tam77b, WD76]. compartmental [WP77]. competing [Gup79b]. complete [Lin79, See78]. complex [Kur78]. Component [HFJ77]. components [GB78, HFJ77]. composite [Loc76, Que78]. compound [Haa78, LA76]. Computation [Amo76]. computational [HM79]. computations [PT76]. Concerning [OB76]. concomitant [Sto77]. conditional [AW76, BK77, Far78, HB78, QS76]. conditions [ZS76]. confidencce [MG79]. Confidence [Cam79a, Fie77b, AS77, ABD77, BK77, Jon77b, LO76, LS79, Mar78a, PW78, Sch76, SL78, UW79]. connected [Mat79, Wyn77]. connectedness [Rus76, Sia77]. conservative [Khu78]. consideration [FPS77]. considerations [KC77]. Consistency [PS76b]. constant [BC79, CB76]. Constrained [McL77, Saw77]. constraints [HR79, Sne79]. contiguity [APR79]. contingency [Boa77, GWD76, KG78, PP78, Rea77]. continuous [UW79]. contours [Bru78]. contrasts [Che79, EHB77]. contributions [Boa77]. convergence [EHB77, Mic76, O’B79]. convex [GW78]. Cornish [Cun77]. Corrections [Tam77a, Sug78]. correlated [Haq78]. Correlation [SB76b, FGS76, HN79, Jam79, JK77, PS77b, Wil78]. cost [Ban78a]. Counterexamples [Ton79, Gup79b]. course [Haa76]. courses [Gun76]. covariance [BT79, Bha77, Bro77b, Fed79, HH79, Hil79, Hui79, KEKL76, NB78, RN79, RY77, Sea79, See78, Sen79]. covariances [PS76a]. covariate [SHW77]. covariates [Ell79]. criteria [CH76]. criterion [DM76, Sug78]. crop [OB76]. cross [Gar78]. cross-product [Gar78]. crossed [Sea79]. crossing [Koo78]. cubic [MK76]. cumulants [Cun77]. Current [Ode76]. curve [KF78]. curves [Zer79].

D. [TB76]. data [ABC79, BT79, BC79, Cam79b, Cli76, Dav78, FH76, Gan78, GK78, HN78, HM79, ISF79, MS77, MH78, MR79, MJB77, OB76, Oma78, Pad77, PS77b, PS79, Sch77, Sug78, WLCC76]. decision [GH77, Lyn78, Tur76]. decisions [Haa78]. decisive [Aro76]. deficient [Bro78c]. definite [MH79]. degree [NB76, NB78, Stu78]. degrees [SW77a, Yam77]. densities [Gup79a]. density [Gra76]. dependence [AW76, Pri76]. dependent [Cha76, GC79, MH78]. derivation [Wil78]. derived [Ban78b]. design [AH78, KC77, Kie78, Kur78, Nor76, STT78, Ski78, TS77, Zer79]. designing [Fed78]. Designs [Stu78, AF76, AT79, CSH77, CI76, DG77, HN78, HA78, Joh77, PF77, PN77a, PN77b, Rus76, SP76a, Sia77, STT78, Sne79, SG77, Wu78]. detecting [HRT77]. detection [BK79]. determinant [Rub76]. determinants [Cun77]. determination [Ram77b]. Developments [Ney79]. deviations [AG78]. diagonal [Bro77b]. dichotomous [SHW77]. difference [MA77]. different [KT78]. dilution [SKT78]. dimension [Hog76]. dimensional [Mar78b]. dimensionality [ST76]. dimensions [Lie77]. Dirichlet [PP78, SP76b]. discrete [Ber78, CK76, GD77, HT78, Lin77, Nel79, PR77, Pee79, VW78]. discrete-time [Pee79]. discriminant [CLB79, Di 76, Di 77, Dip79]. discriminating [Fed78]. Discrimination [GT76a, DH78, GK78, Joh78, McL77]. discussion [Pra76]. dispersion [MH79, Rao78a, SS76]. distributed [Ave79]. Distribution [Gup76, UW79, AS79, ATC78, BEW77, BK78a, BS78, Bru79, Bru78, Cho78, Dah77, Fie77a, FHK76, GM79, GT76b, Gup77a, HG76, Hid78, HF76, HAT78, Hwa76, HT78, Ima76, ID76, JS76, Jon77a, KA79, Kam78, KR79, Koc78, Koz79, Kri76, KC79, Loc76, MCCC76, Mui76, NB76, NS76, PW77, PW79, Par76, Per77, Pet79a, Pet79b, Pet79c, PT76, Res76, Sch77, SW77b, Smi77b, Sub78, SP76b, SL78, Wat79, Wil78]. Distribution-free [UW79, FHK76, Jon77a, Koc78, SW77b]. Distributional [LH79]. distributions [AW76, Amo76, Ber78, BK78b, Che79, CK76, DNV77, FGS76, Gup79a, JK77, Kem78, Kha78, Kor77, LR76, LA76, Leo77, Lin77, MG79, Mat79, Mie79a, Nai76, O’B79, PR77, Pla77, Ram77b, Sha77, Shi78, SS79b, SG76, SO79, TG79b, VW78, WB79, Woo78]. diversity [LH79]. divisible [Joh77]. Double [Spu76]. doubly [MCCC76]. due [Shu76].

EDF [Que78]. Editorial [Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano76p, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano78o, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m]. effect [SG77]. effects [KP78, TS77]. efficiencies [CK76]. Efficiency [DV79, Hwa78, Lep76, PN77b, Sub78]. efficient [Per77]. eigenproblem [Coh77]. elementary [BDL76, Hor76]. ellipticity [Mui76]. emerging [KT78]. Empirical [SO79, SP76b, HK76b, Koz79, Mic76, O’B79, Pet79c, PS79, SS79b, Smi77a, TC76]. English [Dow76]. entry [MS78b]. environmental [Wil76, ZS76]. epidemiology [Bre78]. equal [PS76a]. Equality [Puk77, Day77, Spu76]. equations [Ban78a, PS76b]. equi [Haq78]. equipped [Gen76]. Equivariance [Vin78a]. error [AS79, BH78, Bha77, Bis78, DS78, GM76]. errors [Bha77]. estimability [FR76]. Estimable [BM76, Yam77]. estimate [PC76]. estimated [KEKL76, Woo78]. estimates [AG78, HU77, Hil79, Puk77, SO79, Yam77]. Estimating [Blu77, BS78, Far78, GB78, HB78, McC78, TS77, GW78, GD76, Hil79, McL79, NS76, OB76, Shi78, Wat79, ZS76]. Estimation [AS79, BiK79, BC79, Bro78b, CB76, Dah77, DS77a, HM79, KA79, PW79, Rao79, RW77, Res76, Sen78a, ABC79, Bha77, Bro77b, Bro78c, BM78, CH76, Cli76, DG77, DS77b, DG79, EHB77, GKS77, GM79, Gra76, Gup77a, GW77, HMC79, HK76a, HK76b, KR79, KC77, Law78, Led79, LE77, Leo77, MH79, O’B79, PW77, Pan79, PZW79, Pri76, SKT78, Seh78, Smi77b, SG77, Sub78, VW78, WB79]. estimative [McL79]. estimator [AH79, AGS77, Ban78c, JC76, KC79, RR76, Sha79, Shu76, SR79]. Estimators [GG77, BK79, CK78, Cro77, DL77, DS78, DK79, DKH79, FJ77, GM76, Har78, HR76, Kad79, Kam78, LR76, LW76, MR79, Oma78, Pad77, Pol79, RY77, Rol76, SM76, SM77, Swi76, TC76, Vin78a, WC77, YM76]. evaluating [SMR78c]. Evaluation [Dip79, FJ77, PC76, Rub76]. even [PS76c]. Exact [Che79, MJB79, Pet79a, Ima76, ID76]. Examination [HRT77]. exchangeable [Day77]. existence [SR79]. expansions [Cun77]. expectation [Kou78]. expected [Hut79]. experiment [BSL79, MSVS79, SKT78]. Experimental [Sne79, CI76, TS77]. experimentation [Gab79]. experimenter [Smi79]. experiments [Fed78, HBB^{+}79, HP76, Hui79, Kur78, Mie79b, PN76, PN78, SHW77, SS79b, Smi79, Wei78a, ZS76]. explained [GG77]. Explicit [RY77]. exponential [ABC79, BEW77, BK78a, BY78, DNV77, HF76, HAT78, Kam78, Lin79, Mat79, Per77, Pet79a, Pet79b, Pla77, SG76]. expressing [Hen76]. extended [MJB79, Zer79]. extension [Hsi79, PT76]. extensions [Kri76]. extra [SG77]. extraction [PS76a]. extremal [Fed78]. extreme [LM76, Nai76, Smi77b]. extreme-value [Smi77b].

facto [Ban78b]. Factor [FH76]. factorial [AT79, Ban78a, Bea77, DM79, PN76, PN78, SG77]. factors [Kur78]. failures [PR78]. families [Kie78, TG79b]. family [Pla77]. Farlie [JK77]. Feature [DM76, GW77, GW78, PS76a]. Finite [Rob77, AG78, Kre78, MS78a, Sug78]. Finite-parameter [Rob77]. first [Spu78]. fisher [MJB79, Cun77, SRC78]. fit [Bru78, HMM77, HR76, KF78, Que78, SB76b, TG79a, Wei76]. fitting [Bea77]. Fixed [Jon77b, SL78, PS76c, Tur76]. Fixed-size [Jon77b]. Forecasting [Bro77a]. form [Jam79]. Forms [Mui76, ATC78, GC79, Hoc77]. formulation [HP76]. four [KF78]. fraction [GG77]. fractional [AT79]. free [FHK76, Jon77a, Koc78, SW77b, UW79]. freedom [SW77a]. frequencies [Hut79]. frequency [Rob77]. function [AS79, BM78, CLB79, HF76, Jon77a, KA79, Koz79, LP77, MH78, Mie79c, Pet79c, Rob77, Sha79, SG76, SP76b, Wat79]. functional [RW77]. functions [BM76, Cam79a, FR76, MJB79]. Further [Di 77, Mar78a, NS76, Sug78, SMR78a, DK79, JK77, SD77]. future [Bro77a].

Galton [Ram77b]. Gamma [Ber79, Lin79, Loc76]. Gammaization [GC79]. Gart [BV79]. Gauss [Rao78a, Rao79]. Gaussian [PW79, PS76a]. general [AT79, Bis78, DS76b, GD76, Hoc77, MH79, MH76, Res76]. generalisation [BN76]. generalised [Jam79]. generalization [JS78, SD77]. generalizations [JK77]. Generalized [GM76, AM79, Coh77, Gup76, JK77, Kem78, Lin77, MA77, NB78, Sha79]. generated [AS77, Ney79]. generating [Wu78]. geometric [PT76]. goal [Kad79]. Good [Swi76]. goodness [Bru78, HMM77, HR76, KF78, Que78, SB76b, TG79a]. goodness-of-fit [Bru78, KF78, Que78, SB76b, TG79a]. Graduate [HM76]. graeco [DS77a]. graeco-latin [DS77a]. Graphical [Boa77]. graphics [Gen76]. Gray [AL79]. Grenander [AGS77]. group [Dow76, Joh77, TS77]. grouped [ABC79, BC79, MS77, MR79, Pad77]. growth [Zer79]. Gumbel [JK77]. Guttman [MR78].

half [LE77]. half-sample [LE77]. having [FHK76]. hazard [BC79, CB76, MH78, MJB79, Sha79]. hazards [Bre78]. health [Wil76]. hierarchical [GB78]. Hodges [Pad77]. Hoerl [SMR78c]. Hogg [DV79]. homogeneity [LM76, MB78, TG77]. homogeneous [SRC78]. homoscedasticity [BK78a]. hospital [CM77]. Householder [DS76a]. hypergeometric [GT76a, Hid78]. hypotheses [Kad76, MH76, MD78, Ney76, Pra76, QS76, Rob76]. hypothesis [Kem76, Loc76, Woo78].

I.I.D. [APR79]. ideas [BK77]. identical [Wei78b]. II [JK77, PN78, TB79]. improved [HM79, TB79]. incidence [Gar78]. incomplete [HM79, SB76a]. increments [FM78]. independence [GWD76, MB78, NB78, Que78]. independent [Gup76, Hen76, SW77b, Ski78]. index [Ban78a, Ban78b, LH79, Wil76]. Indifference [ABD77]. Indifference-zone [ABD77]. Inequalities [HG76, PS77b, Rin78, SS79a, Saw77]. inequality [HR79, Ton77]. inference [HN79, Hoc77, Mie79b, Sha77, SG76]. inferences [Bro78a, HB77, Ton79]. Inferential [BEW77]. inflated [Lin77]. influence [Kur78]. Information [KG78, FJ77, FH79, GK78, MJB79, SHW77, Sug78, Swi76, YBW78]. instruction [HM76]. integral [DS76b, QS76]. interacting [Pra79]. interaction [BV79, ZS76]. interactions [KG78]. Interactive [Gen76]. interpolation [Pee79]. interpretation [Ban78a]. interrupted [Hui79]. interval [CH76, LO76, Smi77b, SO79, SL78]. intervals [AS77, ABD77, MG79]. intra [HN79]. intra-class [HN79]. introduction [AH78, GM78, Gup77b, Hog77, Kad76]. Invariance [MS78a, PR77, FR76]. inverse [DM79, Hid78, Jam79, PW79, SD76]. inverses [MA77]. investigating [GD76]. investigation [WLCC76]. involving [HR79, O’B79, Pet79a, Pet79b, SO79]. ipso [Ban78b]. issue [Fed79]. issues [Gab79]. iterative [EHB77, HK76a, Wu78]. iteratively [HW77]. IV [AT79, SL79].

jackknife [LE77, SW77a]. Jackknifing [Kre78, MS78a]. Jonckheere [Ski78].

Kemp [TG79b]. Kennard [SMR78c]. known [Kha78, KR79, Sen78a, Tur76]. Kolmogorov [Sch77, Woo78]. Kruskal [ID76].

Lagrangian [Kor77]. Large [BP77, Stu78]. larger [Gor78]. largest [TB77, TB79]. latin [Add76, DS77a]. lay [RN79]. lay-outs [RN79]. layout [HP76]. layouts [Wyn77]. Least [Nel79, Bea77, Cli76, GKS77, Har77, HW77, Rao78b, Seh78]. least-squares [HW77]. Lehmann [Pad77]. level [FPS77, KP78]. Levels [LA76, DGL79]. life [BT79, Pet79a, Pet79b, Sha77, SS79b]. like [FHK76]. likelihood [Bha77, BM78, DH78, HN79, Kam78, KC79, PC76, RY77, SH78]. limiting [Woo78, Yam77]. limits [Sch76]. Line [ABC79]. linear [Bis78, Bro78a, CK78, Che79, DS76a, DG79, FR76, GW77, GW78, GD76, HT78, Ked77, KEKL76, KG78, LE77, MH79, MH76, Mie79c, Oli78, Oma78, PS76a, Rao78a, RW77, See78, ST76, SM76, SM77, Ton79, YM76, YBW78]. linearity [Adi76a, PR77]. lines [Adi76b]. list [Hut79]. literature [AH78]. Littell [Hsi79]. livelihood [GB78]. living [Ban78a]. Locally [MJB77].

Location [YM76, Haq78, HMC79, KA79, Kam78, Lep76, Lep77, Per77, WC77, Woo78]. location-scale [Haq78]. log [KEKL76]. log-linear [KEKL76]. logarithmic [Gup76]. logistic [AS79, BM78, DH78, MH78]. lognormal [Cro77, PW77]. look [DK79, Har76]. look-up [Har76]. loss [Mie79c]. lower [TB79].

m.s [Che78]. mail [HA78]. Mann [Ima76, FHK76]. Marginal [MB78, HN79]. Markoff [Rao78a, Rao79]. Markov [BiK79, CK78, Hui79, WP77]. matched [BS76, WD76]. Mathematical [ATC78, Ney79]. matrices [Bro77b, MA77, RY77]. Matrix [Gra76, Bha77, Hil79, Hwa76, MH79, Rao78a, See78, SS76]. max [McC78]. maximin [Hwa76]. Maximum [Bha77, Kam78, SH78, BM78, GB78, KC79, PC76, RY77]. maximum-likelihood [PC76]. mean [BH78, Ban78c, BS78, BP77, DS78, GM79, GM76, HRT77, KR79, MG79, Sen78a, Shi78, SR79, TB77, TB79, Wat79]. meaningful [Ell79]. means [Ave79, Che79, Cro77, DV79, FM78, Hua76, Kha78, San76, Spu76]. measure [AL79]. mechanisms [KT78]. median [SL78]. meteorology [Ney79]. method [HM79, McC78, Ram77b, AM79]. Methods [GW78, AS77, BM78, CM77, FPS77, McL79, OB76, PR78, SMR78c, VW78]. Minimax [AH79, Ber79, TB77, TB79]. Minimum [Cro77, Gup77a, SM76, AS79, BH78, DS78, GK78, KC77, SM77]. Minium [Seh78]. MINQUE [Bro77b, GB78, RC78, SP76a]. misclassification [Dip79]. missing [DS77a, HN78, PS77b, WLCC76]. mixed [Bro78a, HH79, SH78, Ton79]. mixture [PC76, Scl77, Sne79, TC76]. mode [AGS77]. model [Bha77, BM76, Bis78, BC79, Bre78, CB76, DG79, DS78, Haq78, Hui79, KEKL76, MH76, Oli78, Rao78a, Rao79, RW77, San76, Sch76, See78, SH78, Ste79, Tam77a, Tam77b]. models [And78, AG78, BT79, Bro78a, Bro78c, Coh77, DG78, FR76, GT76a, HH79, KF78, Led79, Mar78a, MH79, McC78, MD78, PP78, Scl77, Sea79, Sen79, SM76, SM77, TB76, Ton79, WP77]. Modification [PS76c, BSL79, HBB^{+}79, Mie79b, Ney79, RF79]. modifications [RC78]. modified [Gup77a, HP76, KC79, MR78, O’B79, WB79, YM76]. modular [HM76]. moment [Ber78, HR76]. Moments [Aro76, Sco76, GT76b, Hid78, KC79, Ram77b]. monotonic [HRT77]. monotonicity [AW76, Gor78]. Monte [DL77, WC77, WLCC76]. Morgenstern [JK77]. most [MJB77, SKT78]. Mosteller [JS78]. moving [McC78, Pee79]. MRPP [Mie79a]. MSE [FJ77]. Multi [MBJ76, Mar78b]. multi-dimensional [Mar78b]. Multi-response [MBJ76]. multicollinear [Oma78]. Multicollinearity [FH79]. multicomponent [DG78, Sne79]. Multidimensional [AF76, CSH77]. multinomial [Leo77]. multinormal [GT76b]. Multiple [BS76, Tam77a, Tam77b, Tur76, Agr77, AM79, GH77, Haa78, HL76, KEKL76, KR77, MR78, MJB77, SB76a, Wil78]. multiple-censored [MJB77]. multivariable [ST76]. multivariate [Bha77, BM78, Coh77, DS76b, Fie77a, HG76, HMM77, Hsi79, PW78, Wat79].

N [Dra76]. natural [Ney76]. Negative [KC79, Bof78, HF76, Sch76]. nested [HM79]. noisy [Cam79b]. Non [PZW79, Bof78, CLB79, Mat79, Mie79a, RN79]. non-negative [Bof78]. non-normality [CLB79, Mie79a]. non-null [Mat79]. non-orthogonal [RN79]. Non-parametric [PZW79]. noncentral [Amo76, Kri76, MCCC76, NS76]. noncentral- [MCCC76]. nonfull [GD76]. nonlinear [BM76, MD78]. nonnegative [MA77]. nonnormal [Ban78c]. Nonparametric [HK76b, MS78b, Sen79, CK78, Dur76, Gov76, Hog76, Lep77, Mie79b, Smi77a]. nonsingular [Wu78]. nonsymbolic [Haa76]. Nonuniqueness [Har77]. normal [ATC78, Ban78c, BS78, Bru79, Bru78, DK79, Fie77a, FM78, HG76, Hua76, Kha78, KR79, LA76, Mar78a, MG79, NB76, Pan79, PW78, PS76b, San76, Shi78, TB77, TB79, TS77, Wat79]. normality [BP77, CLB79, GS78, HMM77, Ked77, Mie79a, Pet79c, See78]. normally [Ave79]. normals [GW78]. norms [HFJ77]. Note [Pla77, Adi76a, AL79, BN76, Bru79, Che78, Dev77, Hen76, Kha78, PS76a, Pri76, Rao78b, Sch76, Shi78, SM77, Vin78a]. null [FHK76, Mat79, Mie79a, Woo78]. null-hypothesis [Woo78]. number [Aro76, KA79, Mic76, Oli78, PF77, Pla77, SKT78]. numbers [HH79]. Numerical [Nak79, PC76].

O.L.S. [MH79]. observations [Bof78, Cha76, DS77a, DG79, LA76]. obtained [ISF79]. one [BS78, GG77, HF76, HP76, Jon77a, Sch77, SG77, Tam77a, Tam77b]. one-sample [HF76, Jon77a]. one-sided [Sch77]. one-way [GG77, HP76, Tam77a, Tam77b]. onesided [Ave79]. Operational [Kad79]. Optimal [DS76a, Kur78, BK78b, Cam79b, CSH77, HR79, QS76, San76, STT78, Wu78]. optimality [AH78, Che78, CH76, Rus76]. optimum [Dip79, PS76a]. order [BV79, KA79, MS77, McC78, Mic76, Rub76, YM76]. ordered [Adi76b, SW77b]. ordinary [BY78]. origin [HRT77]. orthogonal [RN79]. other [LW76, Ton77]. Outlier [BK79, Tik77]. Outlier-detection [BK79]. outliers [GD76, Haw77, MR78, Tik75]. outs [RN79]. overview [AH78].

P [ABD77]. pair [Ban78a]. paired [Bea77, EHB77]. Pairwise [WD76]. paradox [Gup79c]. parallelism [Adi76b]. parameter [Adi76a, APR79, Cam79b, HK76a, KR79, LO76, Led79, MR79, NS76, Nel79, PW77, PW79, Par76, PZW79, Per77, Pet79a, Pet79b, Pra79, Rob77, Smi77b]. parameters [Day77, DS77a, HMC79, KA79, Kam78, KEKL76, LM76, Lep77, Rao79, Res76, WB79, Woo78, Yam77]. parametric [FR76, Mie79b, PZW79]. partial [AL79, Day77]. Partially [PN77b, PF77]. Partitioning [JS76, Rea77, FM78]. Pascal [Sub78]. past [Ode76]. patterned [RY77]. patterns [KR77]. PCS [TB79]. Pearson [SL79]. survivability [DKH79]. percentile [DK79]. periodic [Cam79a]. permutation [MBJ76, Mie79b]. Perng [Hsi79]. phase [HB77]. phenomena [Ney76]. Physically [Ell79]. piece [BC79, CB76]. piece-wise [BC79, CB76]. planning [HN78]. play [Bof78, PS76c]. play-the-winner [Bof78, PS76c]. point [GM79]. Poisson [Dah77, Gar78, Gbu79, Kem78, Kor77, O’B79, Sub78, WB79]. policy [BK78b]. polynomial [NB76, NB78, Stu78]. Population [Scl77, ABC79, AG78, Ban78c, Blu77, BP77, HAT78, KW77, Koo78, Loc76, MS78a, Pan79, Sen78a]. populations [BM77, Day77, FM78, HMC79, Kre78, Lin79, Mat79, PS77a, PS76a, Tur76, Wei78b]. Positive [AW76, Hid78, Hoc77, MH79]. posterior [BP77, McL79]. potency [RR76]. potential [HN78]. power [BY78, Gor78, Gup77a, HF76, Jon77a, KA79, KC79, Lep76, Par76]. powerful [MJB77]. practice [Law78]. precision [PW78, Shu76]. predicted [AS79]. Predicting [CM77]. prediction [Haq78, Kad79, Lin79, Pee79]. predictive [McL79]. predictor [Ram77a]. Preface [Fed79, WO79, Sen78b]. preliminary [BS78]. prescribed [PW78]. present [DG79]. principal [HFJ77]. principle [RC78]. principles [MS78a]. prior [Ban78c, FJ77, Swi76, YBW78]. priori [FH79, MBJ76, PS76a]. priors [SP76b]. probabilities [BiK79, DS76b, HG76, Kem78, McL79, PS76a, Rub76, SL79]. Probability [PT76, Bru78, Dip79, HK76b, Koo78, Ney79, QS76, Rin78, Ton77]. probability-inequalities [Rin78]. probable [SKT78]. problem [DH78, Koc78, Lep76, Lin79, O’B79, Par76, Pet79a, Pet79b, Shi78, SRC78, WLCC76]. problems [Fed78, Fie77b, GH77, Hua76, HR79, Kie78, Nai77, PS79, Rol76, SS79b, Ton77]. procedure [Bis78, CSH77, DNV77, GD77, MD78, MK79, PS76c, San76, SP76a, TB77, TB79]. Procedures [GD76, BEW77, BM77, CI76, Day77, Don77, Gup77b, HL76, Hog77, Hua76, Law78, MBJ76, Mie79b, Muk79, PS77a, Rin78, Sen78b, Wu78, Yue77]. process [BY78, Ram77b, SP76b]. processes [Nel79]. product [ATC78, Gar78, Hen76]. programming [Fie77b]. progressive [MS77, MS78b]. Progressively [SS79b, Dav78]. proper [RW77]. Properties [LE77, Sia77, Ber78, GM76, Gup79a, HU77, LH79, Mar78b, Que78, SS76]. property [AH79]. proportion [TC76]. proportional [Bre78]. proportions [BS76, GW77, GW78, Khu78, PC76, WD76]. protection [FPS77]. providing [Ban78a].

quadratic [CLB79, GC79, Hoc77, MK76]. qualitative [Kur78]. Quantal [SD76, DM79]. quantiles [AS79, Cha76, PW78]. quasi [Kor77]. quasi-binomial [Kor77]. question [Pra76]. questions [Dra76].

Random [And78, Hwa76, ISF79, KP78, MS78b, Pla77, Saw77, Sco76]. Randomization [Zer79]. randomized [Dev77, Dra76, FPS77, Ski78, Zer79]. randomness [Lie77]. Rane [FHK76]. Rane-like [FHK76]. range [HL76]. ranges [Don77]. Rank [MS77, PS77b, Bro78c, Gor78, GD76, HU77, Hwa76, Hwa78, HT78, Ima76, Ked77, MJB77]. Ranked [Sto77]. ranking [ABD77, Gup77b, Ton77]. rankings [SB76a]. ranks [Agr77, CI76, MH76]. rare [PR78]. rate [BC79, CB76, MJB79, Sha79]. Rates [O’B79, APR79, Gar78]. ratio [Cro77, Cun77, DH78, DS77b, HAT78, JC76]. ratios [Bro78a, FM78, Gar78, Ton79]. Reassessing [Gun76]. recapture [KEKL76]. recidivism [BT79]. recognition [Ano77p]. rectangle [Rub76]. rectangular [Jon77b]. Recurrence [GT76b]. recursive [Led79, Ram77b]. reduction [ST76]. Referee [Ano77p]. reference [HP76]. references [Hut79]. regions [Cam79a, Jon77b]. regression [Adi76a, Adi76b, AG78, BH78, Bha77, CM77, DL77, DM79, DS77b, DS78, Far78, GM76, HB78, Har77, Har78, HRT77, HFJ77, Hil79, HK76a, HB77, HW77, JK77, Kad79, LW76, Led79, LP77, MS78b, MH78, Oli78, Oma78, PR77, RW77, Rol76, SD76, SS76, Stu78, SMR78c, YBW78]. regressions [SR79]. rejection [DGL79, Spu78]. Rejoinder [Tik77, Vin78b]. related [HB77, Law78, PR77, Rin78, Ton77]. relation [HRT77]. Relations [Yam77, GT76b]. relationship [GM78, Oli78]. relative [DKH79, RR76]. relatively [Mic76]. reliability [DG78, DKH79, Gup79a, PW77, SG76]. reliability/survivability [DKH79]. remarks [NS76]. remote [OB76]. renewal [BY78, Gup79a]. repeated [Sen79]. Reply [SMR78b, SMR78a]. representation [DS76b]. requirements [PW78]. research [Hut79]. residuals [BK79, HRT77]. Resolution [AT79, SG77]. respondent [FPS77]. response [Bea77, Dev77, DM79, FPS77, Haq78, MH78, MK76, MBJ76, Rob77, SD76, SHW77, Zer79]. response-time [MH78]. responses [Dra76, Gbu79]. restricted [GB78, Mar78a]. result [Mie79b, Ton79]. results [Mar78a, MSVS79, Smi77b]. revisited [Sub78]. reweighted [HW77]. Ridge [HK76a, Law78, BH78, Bro78c, FJ77, Kad79, LW76, Puk77, SMR78c, Swi76, Vin78a]. ridge-type [Puk77]. ridits [Lyn78]. risk [BM78, Gup79b, SM76, SM77]. Robust [DL77, HW77, Len77, SK77, BK79, HMC79, Har78, Hil79, Hog77, Pol79, TG77, WC77]. Robustness [Ban78c, HU77, Yue77, Ber79, DV79, HA78, HF76]. roles [Ode76]. row [Rus76]. row-column [Rus76]. rule [Har76, MR78]. rules [Ber79, Tur76, Wei78a]. Ryan [Shu76].

Sample [APR79, Khu78, PW78, Aro76, Bis78, BP77, Cha76, CK76, Dav78, Don77, Gor78, HF76, Hua76, HR79, Jon77a, Kam78, Koc78, LO76, LA76, LE77, Man78, McL77, Oli78, Par76, PS79, PS76c, RF79, Spu76, Wei76]. Sample-size [PW78]. samples [HAT78, KR77, Lin79, MG79, MJB79, TS77, Wei78b]. sampling [Blu77, Bof78, BK78b, BS78, Gup79c, ISF79, MS78a, SB76b, Sto77]. savage [SD77]. Scale [LR76, Dur76, Haq78, HMC79, KA79, Kam78, LO76, LM76, Lep76, Lep77, Per77, Woo78]. scales [Ski78, TG77]. scaling [Mar78b]. Scheffé [AS77, AM79]. scheme [BS78, CK78, EHB77, PF77]. schemes [DS77b]. school [BDL76, Hor76]. schools [Dow76]. scores [BY78]. screening [TB77, TB79]. search [SG77]. second [BV79, NB76, NB78]. Secondary [KC77]. seeding [MSVS79, Smi79]. seemingly [SR79]. selected [KA79]. Selecting [KP78, Ram77a, DNV77, KW77, San76, TB77, TB79]. Selection [Bof78, Gup77b, Joh78, ABD77, BM77, Don77, GD77, GW77, GW78, HFJ77, HP76, Hua76, Mie79c, Muk79, Nai77, PS77a, Rin78, Ton77]. semi [WP77]. semi-Markov [WP77]. semiregular [Joh77]. sensing [OB76]. sensitive [Dra76]. sensitivity [TC76]. separate [QS76]. sequences [Bro77a, HB77, LS79]. Sequential [BM77, GM79, Aro76, DS77b, GS78, LS79, Nor76, RF79, Sen78b, SHW77, SL78, Wei78a, Yue77]. Serial [FGS76, SKT78]. Series [Joh77, CM77, Gup76, Gup77a, Hui79, KC79, Led79, Par76]. service [Gun76]. set [Sto77]. sets [Fie77b, SB76a]. setting [BK77]. several [Khu78, Leo77, OB76, Tur76, Wei78b]. shape [Cam79b]. shifting [HB77]. shrinkage [Pan79, Rol76]. sided [Sch77]. sign [Spu78]. Significance [Ave79, BS78, Kad76, Kem76, LA76, Pra76, Rob76, Sen79]. signs [BK79]. similar [SS79a]. similarity [Que78]. simple [DH78, HU77, MS78b, Wil78]. simplest [BK77]. Simpson [LH79]. simulation [LW76]. Simultaneous [Bro78a, CH76, Hoc77, Leo77, Mar78a, Ton79]. singly [Bru78]. Singular [DG77, MH79, NB76, Rao78a, Rao79, See78]. situation [GD76]. situation-nonfull [GD76]. size [APR79, ABC79, Blu77, Gup79c, Jon77b, Kem78, Khu78, Oli78, PW78, PS76c]. sizes [Hua76]. skewness [DQ78]. slippage [Koc78]. slope [RW77]. slopes [Pet79a]. Slustep [Kem78]. Small [MG79, Hut79, Man78, Mic76]. small-sample [Man78]. smallest [Bro78b]. Smiirnov [Woo78]. Smirnov [Sch77]. Smooth [VW78]. smoothing [Cam79b]. Some [BSL79, Cha76, Fed78, FHK76, Gab79, GH77, Gup79b, HR79, Mar78b, Mie79b, Muk79, Nai77, PR78, Sia77, Smi77b, TG79b, Wei78b, Wu78, Bro78a, CK78, CI76, Gan78, Gup77a, Hua76, HT78, JK77, LA76, Lin77, MS78a, Oma78, PR77, Sen79, Ton79, Yue77]. space [Ode76, STT78]. special [Ban78a, Fed79, HP76]. species [Pra79]. specification [Far78, HB78]. sphere [FGS76]. sphericity [Mui76]. spline [Pee79]. splines [Len77]. spurious [DG79]. square [NS76, PP78, Rea77]. squared [BH78, DS78, GM76, Hut79, Kri76]. squares [Add76, Bea77, Cli76, DS77a, HW77, Nel79, Rao78b, Seh78]. Stability [Pra79]. stable [Jon77a, LR76]. stage [BS78, DNV77, HL76, Muk79, Rin78, San76, Spu78, TB77, TB79]. staggering [MS78b]. standard [Tur76]. standardization [Vin78a]. state [BiK79]. stations [ZS76]. Statistic [Tik77, Adi76a, Boa77, Haw77, Ima76, ID76, LO76, McL77, Mui76, NB76, Sch77, See78, ST76, Tik75, WLCC76]. Statistical [Sha77, BSL79, Coh77, Gab79, Ney76, PR78, Sen78b]. statistician [Ode76]. Statistics [BDL76, Dow76, DQ78, FHK76, GS78, Gov76, Gun76, Haa76, Hor76, KA79, Ked77, Kre78, MS78a, Mat79, Mic76, Mie79a, NB78, Ney79, Rub76, SS79a, SW77b, Ski78, SB76b, Smi79, Woo78, Yam77, YM76, GH78]. status [Wil76]. Steck [Rub76]. steepest [MK79]. Stein [AH79, Bis78, Shi78]. stepwise [DGL79, GD77]. stochastic [FJ77, Nel79, Pra79, ST76, WP77]. stochastically [Gor78]. stratified [ISF79]. strength [DG78]. stress [DG78]. stress-strength [DG78]. structural [TB76]. structure [KEKL76]. structures [GB78]. Student [HAT78, Res76, SS79a]. Studentized [HL76, McL77]. students [Hor76]. studies [Ney76, Ney79]. study [LW76, TC76, WC77]. subclass [HH79]. subject [KR77, Pra79]. subpopulation [JC76]. Subset [Don77, PS77a, HP76, Mie79c, Nai77]. successive [FM78]. sufficient [See78]. Sukhatme [Hsi79]. Sukhtmes [LO76]. sum [Gup76, Ima76]. sums [Pla77]. superior [PS77a]. suprema [Sco76]. surface [Bea77]. survey [BT79, Dur76, FPS77, Gov76]. surveys [HR79]. survival [Gup79a, KF78, PR78]. Suspected [Tik77, Haw77, Tik75]. symmetric [Coh77, HMC79, LR76, UW79]. symmetry [DQ78, Koz79, MB78, Smi77a]. system [ST76]. Systematic [Add76]. systems [Sne79].

table [Har76]. Tables [KT78, Boa77, GWD76, KG78, PP78, Ray76, Rea77, TB79]. tail [HG76, SL79, TG79a]. Tasmania [MSVS79]. Teaching [Hor76, Rea77]. technique [Dev77, DS76b]. techniques [LE77]. terminal [Gen76]. terminal-equipped [Gen76]. termination [Aro76]. test [BV79, BS78, Bru78, BY78, Dav78, DH78, DGL79, GWD76, HRT77, HMM77, HF76, Hsi79, Hut79, Hwa78, Ima76, ID76, Jon77a, JS78, Kou77, Koz79, KP78, Loc76, LS79, LP77, Mat79, MD78, Per77, Ray76, RF79, SD77, Smi77a, TG77]. Testing [Adi76b, Kou78, Pet79c, SW77b, Tik77, Wil76, Day77, Gor78, Haw77, Kha78, MS78b, Man78, Pet79a, Pet79b, SS79b, SP76b, TG79a, Tik75, TS77, ZS76]. Tests [LM76, Lie77, MH76, Ney76, Aro76, Ave79, BK79, CK76, Dur76, Gor78, Hog76, HR76, HT78, Kad76, Kem76, Koc78, LA76, Lep77, LS78, MS77, MJB77, Nel79, Pra76, QS76, Que78, Rob76, Sen79, SS79b, Spu76, Spu78, Wei76, Wei78b]. their [Ber78, Gup79a, NB78, Ney76]. theorem [BN76, FR76, MA77, TB76]. theoretic [Lyn78]. theory [Dra76, Law78, Rao78b]. Three [RC78, PW79, Wyn77]. three-parameter [PW79]. three-way [Wyn77]. thresholds [Ave79]. Tiao [MR78]. ties [HT78]. Time [Sen78b, Aro76, CM77, GS78, Gup79c, Hui79, Led79, MH78, Pee79]. Time-sequential [Sen78b, GS78]. tinder [HB78]. tolerance [Man78]. tool [FH76]. total [DG77, JC76]. transect [ABC79]. transformations [DS76a, QS76]. transition [BiK79]. Treatment [EHB77, HT78, Wei78a]. treatments [ZS76]. trend [BY78]. trials [Nor76, Pri76]. trimmed [DV79]. true [ABD77, Ban78a, Ban78b]. truncated [BEW77, Blu77, Bru78, Dah77, GT76b, KR79, Sch77]. truncation [Par76]. Two [HL76, Par76, Spu78, SMR78c, TB77, Wei76, ATC78, BiK79, BM77, Bis78, Boa77, BS78, CK76, Cro77, Dav78, Day77, DNV77, Dra76, FPS77, Gor78, HB77, HP76, Koc78, LO76, LA76, Lie77, MG79, MA77, Muk79, NB78, PW77, Pet79a, Pet79b, PS79, Pra79, Puk77, RF79, Rin78, San76, Sia77, Smi77b, Spu76, TB79, WB79, ZS76]. two-parameter [PW77, Smi77b]. two-phase [HB77]. Two-sample [Wei76, Bis78, Gor78, Koc78, LO76, LA76, PS79]. Two-stage [TB77, BS78, DNV77, Muk79, Rin78, TB79]. two-state [BiK79]. two-way [Boa77, HP76, Sia77]. type [HR76, Puk77, RF79, SL79, SB76b, Woo78, TS77].

U [DQ78]. U-statistics [DQ78]. unbiased [Cro77, Gup77a]. unbiasedness [Har78]. uncertain [BK78b]. undergraduate [HM76]. underlying [HT78]. Unequal [Tam77a, Bis78, HH79, Hua76, KR77, RF79, Ski78, Tam77b]. unfair [Joh78]. unified [Rao78b]. Uniform [Mic76, Rub76, SO79]. uniformity [LS78]. univariate [GD76]. unknown [Bha77, Par76, SG77]. unrelated [SR79]. unspecified [GM79]. upon [MR79]. Use [TG79a, Hoc77, Kad76, Kem76, Ney76, Pra76, Rob76, SHW77, YBW78]. used [Kou77, Kou78, Oma78]. useful [HN78, ST76]. using [Bof78, BS78, CM77, CI76, DS76a, DM79, GB78, GW78, HW77, OB76, Pet79c, RF79]. Utilizing [DS77b].

V [SG77]. validity [Hut79]. value [FH79, LM76, Nai76, Smi77b]. values [Bro77a, Bro78b, GKS77, Har77]. variable [Bha77, GD77]. variables [ATC78, Ave79, DS77b, Gup76, Hen76, Oli78, Pla77, Saw77, Sco76, Sto77]. variance [AL79, Bro78a, Cro77, Gbu79, GB78, GG77, Gup77a, GH77, HRT77, LE77, Pan79, SK77, Ton79]. Variances [Tam77a, Bis78, Che79, FM78, MG79, RF79, Tam77b]. variate [Ram77a]. variates [Cro77, Gar78, NB76]. variation [KR79, Pra79, Sen78a, SRC78]. variations [Kha78]. via [DL77]. view [Smi79]. visibility [And78].

Waiting [Gup79c]. Wald [Ban78b]. Wallis [ID76]. Watson [Ram77b]. way [Boa77, GG77, HP76, Sea79, Sia77, Tam77a, Tam77b, Wyn77]. weak [BK78a]. weather [BSL79, Gab79, HBB^{+}79, Mie79b, Ney79]. Weibull [Man78, Smi77b]. weighing [DG77]. weight [DG77]. Weighted [Bea77, SS79b]. Welch [RF79]. where [MH78]. whether [Wei78b]. which [SG77]. Whitney [FHK76, Ima76]. width [SL78]. Wilcoxon [Dav78, FHK76, Hwa78, Ima76]. Wilks [Hen76]. Williams [AL79]. winner [Bof78, PS76c]. wise [BC79, CB76]. Wishartness [GC79]. withdrawal [MS78b]. without [Cun77, Joh78].

zone [ABD77].

References

Anderson:1979:LTE

[ABC79] D. R. Anderson, K. P. Burnham, and B. R. Crain. Line transect estimation of population size: the exponential case with grouped data. Communications in Statistics: Theory and Methods, 8(5): 487–507, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1977:IZR

[ABD77] Peter O. Anderson, Thomas A. Bishop, and Edward J. Dudewicz. Indifference-zone ranking and selection: confidence intervals for true achieved P(CD). Communications in Statistics: Theory and Methods, 6(11):1121–1132, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Addelman:1976:SLS

[Add76] Sidney Addelman. Systematic latin squares. Communications in Statistics: Theory and Methods, 5(1):27–44, 1976. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adichie:1976:NAL

[Adi76a] J. N. Adichie. A note on asymptotic linearity of a bank statistic in regression parameter. Communications in Statistics: Theory and Methods, 5(2):171–181, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adichie:1976:TPR

[Adi76b] J. N. Adichie. Testing parallelism of regression lines against ordered alternatives. Communications in Statistics: Theory and Methods, 5(11):985–997, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1976:MBD

[AF76] D. A. Anderson and W. T. Federer. Multidimensional balanced designs. Communications in Statistics: Theory and Methods, 5(13): 1193–1204, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:1978:BDE

[AG78] Barry C. Arnold and Richard A. Groeneveld. Bounds on deviations of estimates arising in finite population regression models. Communications in Statistics: Theory and Methods, 7(12):1173–1179, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agresti:1977:CMA

[Agr77] Alan Agresti. A coefficient of multiple association based on ranks. Communications in Statistics: Theory and Methods, 6(14):1341– 1359, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andriano:1977:ABG

[AGS77] Kim Andriano, James E. Gentle, and V. A. Sposito. On the asymptotic bias of Grenander’s mode estimator. Communications in Statistics: Theory and Methods, 6(8):773–776, 1977. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ash:1978:IDO

[AH78] A. Ash and A. Hedayat. An introduction to design optimality with an overview of the literature. Communications in Statistics: Theory and Methods, 7(14):1295–1325, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1979:MPS

[AH79] Khursheed Alam and James S. Hawkes. Minimax property of Stein’s estimator. Communications in Statistics: Theory and Methods, 8 (6):581–590, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1979:NAV

[AL79] Robert J. Anderson and J. Richard Landis. A note on the asymptotic variance of the Gray and Williams measure of partial association. Communications in Statistics: Theory and Methods, 8(13): 1303–1314, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Altschul:1979:GSM

[AM79] Roberto E. Altschul and Richard Marcuson. A generalized Scheffé method of multiple comparisons. Communications in Statistics: Theory and Methods, 8(3):271–281, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amos:1976:CCN

[Amo76] D. E. Amos. Computation of the central and noncentral F distributions. Communications in Statistics: Theory and Methods, 5(3):

261–281, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1978:RVM

[And78] C. L. Anderson. Random visibility models. Communications in Statistics: Theory and Methods, 7(8):725–731, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:E

[Ano76a] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 5(12):v, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBa

[Ano76b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(1):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBb

[Ano76c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(2):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBc

[Ano76d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(3):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBd

[Ano76e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(4):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBe

[Ano76f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(5):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBf

[Ano76g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(6):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBg

[Ano76h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(7):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBh

[Ano76i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(8):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBi

[Ano76j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(9):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBj

[Ano76k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(10):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBk

[Ano76l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(11):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBl

[Ano76m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(12):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBm

[Ano76n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(13):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBn

[Ano76o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(14):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1976:EBo

[Ano76p] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 5(15):ebi, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBa

[Ano77a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(1):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBb

[Ano77b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(2):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBc

[Ano77c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(3):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBd

[Ano77d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(4):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBe

[Ano77e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(5):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBf

[Ano77f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(6):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBg

[Ano77g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(7):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBh

[Ano77h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(8):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBi

[Ano77i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(9):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBj

[Ano77j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(10):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBk

[Ano77k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(11):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBl

[Ano77l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(12):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBm

[Ano77m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(13):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBn

[Ano77n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(14):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:EBo

[Ano77o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 6(15):ebi, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1977:RR

[Ano77p] Anonymous. Referee recognition. Communications in Statistics: Theory and Methods, 6(15):1531–1533, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBa

[Ano78a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(1):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBb

[Ano78b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(2):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBc

[Ano78c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(3):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBd

[Ano78d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(4):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBe

[Ano78e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(5):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBf

[Ano78f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(6):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBg

[Ano78g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(7):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBh

[Ano78h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(8):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBi

[Ano78i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(9):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBj

[Ano78j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(10):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBk

[Ano78k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(11):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBl

[Ano78l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(12):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBm

[Ano78m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(13):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBn

[Ano78n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(14):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1978:EBo

[Ano78o] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 7(15):ebi, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBa

[Ano79a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(1):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBb

[Ano79b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(2):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBc

[Ano79c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(3):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBd

[Ano79d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(4):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBe

[Ano79e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(5):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBf

[Ano79f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(6):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBg

[Ano79g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(7):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBh

[Ano79h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(8):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBi

[Ano79i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(9):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBj

[Ano79j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(11):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBk

[Ano79k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(13):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBl

[Ano79l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(14):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1979:EBm

[Ano79m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 8(15):ebi, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akritas:1979:SSP

[APR79] M. G. Akritas, M. L. Puri, and G. G. Roussas. Sample size, parameter rates and contiguity: the I.I.D. case. Communications in Statistics: Theory and Methods, 8(1):71–83, 1979. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aroian:1976:MDS

[Aro76] Leo A. Aroian. Moments of the decisive sample number and of the average time to termination of sequential tests. Communications in Statistics: Theory and Methods, 5(11):1041–1045, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alt:1977:CCI

[AS77] Frank Alt and Carl Spruill. A comparison of confidence intervals generated by the Scheffé and Bonferroni methods. Communications in Statistics: Theory and Methods, 6(15):1503–1510, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1979:ELQ

[AS79] J. E. Angus and R. E. Schafer. Estimation of logistic quantiles with minimum error in the predicted distribution function. Communications in Statistics: Theory and Methods, 8(13):1271–1284, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1979:RIF

[AT79] Donald A. Anderson and Ann M. Thomas. Resolution IV fractional factorial designs for the general asymmetric factorial. Communications in Statistics: Theory and Methods, 8(9):931–943, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aroian:1978:MFD

[ATC78] Leo A. Aroian, Vidya S. Taneja, and Larry W. Cornwell. Mathematical forms of the distribution of the product of two normal variables. Communications in Statistics: Theory and Methods, 7(2):165–172, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Avenhaus:1979:STO

[Ave79] Rudolf Avenhaus. Significance thresholds of onesided tests for means of bivariate normally distributed variables. Communications in Statistics: Theory and Methods, 8(3):223–230, 1979. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1976:PDM

[AW76] Khursheed Alam and K. T. Wallenius. Positive dependence and monotonicity in conditional distributions. Communications in Statistics: Theory and Methods, 5(6):525–534, 1976. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1978:FAP

[Ban78a] K. S. Banerjee. On the factorial approach providing the true index of cost of living: an interpretation of a special pair of equations. Communications in Statistics: Theory and Methods, 7(9):851–856, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1978:WTI

[Ban78b] Kali S. Banerjee. Wald’s “true index” derived ipso facto. Communications in Statistics: Theory and Methods, 7(8):791–798, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:1978:RBE

[Ban78c] Ashok K. Bansal. Robustness of a Bayes estimator for the mean of a normal population with nonnormal prior. Communications in Statistics: Theory and Methods, 7(5):453–460, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boardman:1979:EPW

[BC79] Thomas J. Boardman and Robert E. Colvert. Estimation in the piece-wise constant hazard rate model — when the data are grouped. Communications in Statistics: Theory and Methods, 8(5):417–432, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beck:1976:SES

[BDL76] Betty Beck, Lorraine Denby, and James M. Landwehr. Statistics in the elementary school. Communications in Statistics: Theory and Methods, 5(10):883–894, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beaver:1977:WLS

[Bea77] Robert J. Beaver. Weighted least squares response surface fitting in factorial paired comparisons. Communications in Statistics: Theory and Methods, 6(13):1275–1287, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berg:1978:CCD

[Ber78] Sven Berg. Characterization of a class of discrete distributions by properties of their moment distributions. Communications in Statistics: Theory and Methods, 7(8):785–789, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berger:1979:GMR

[Ber79] Roger L. Berger. Gamma minimax robustness of Bayes rules. Communications in Statistics: Theory and Methods, 8(6):543–560, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bain:1977:IPT

[BEW77] Lee J. Bain, Max Engelhardt, and F. T. Wright. Inferential procedures for the truncated exponential distribution. Communications in Statistics: Theory and Methods, 6(2):103–111, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baldwin:1978:BMM

[BH78] Kent F. Baldwin and Arthur E. Hoerl. Bounds on minimum mean squared error in ridge regression. Communications in Statistics: Theory and Methods, 7(13):1209–1232, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhargava:1977:MLE

[Bha77] Anil K. Bhargava. Maximum likelihood estimation in a multivariate ‘errors in variables’ regression model with unknown error covariance matrix. Communications in Statistics: Theory and Methods, 6(7): 587–601, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bai:1979:ETP

[BiK79] Do Sun Bai and Seong in Kim. Estimation of transition probabilities in a two-state Markov chain. Communications in Statistics: Theory and Methods, 8(6):591–599, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bishop:1978:STS

[Bis78] Thomas A. Bishop. A Stein two-sample procedure for the general linear model with unequal error variances. Communications in Statistics: Theory and Methods, 7(5):495–507, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brownie:1977:ICC

[BK77] C. Brownie and J. Kiefer. The ideas of conditional confidence in the simplest setting. Communications in Statistics: Theory and Methods, 6(8):691–751, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beq:1978:CED

[BK78a] M. I. Beq and S. N. U. A. Kirmani. Characterization of the exponential distribution by a weak homoscedasticity. Communications in Statistics: Theory and Methods, 7(3):307–310, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowerman:1978:AOP

[BK78b] Bruce L. Bowerman and Anne B. Koehler. Ax optimal policy for sampling from uncertain distributions. Communications in Statistics: Theory and Methods, 7(11):1041–1051, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1979:ODT

[BK79] B. M. Brown and D. G. Kildea. Outlier-detection tests and robust estimators based on signs of residuals. Communications in Statistics: Theory and Methods, 8(3):257–269, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blumenthal:1977:EPS

[Blu77] Saul Blumenthal. Estimating population size with truncated sampling. Communications in Statistics: Theory and Methods, 6(4): 297–308, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bird:1976:EFN

[BM76] Howard A. Bird and George A. Milliken. Estimable functions in the nonlinear model. Communications in Statistics: Theory and Methods, 5(11):999–1012, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barron:1977:SSP

[BM77] Austin M. Barron and Eugene Mignogna. Sequential selection procedures for choosing the better of two binomial populations. Communications in Statistics: Theory and Methods, 6(6):525–552, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Byth:1978:BAM

[BM78] K. Byth and G. J. McLachlan. The biases associated with maximum likelihood methods of estimation of the multivariate logistic risk function. Communications in Statistics: Theory and Methods, 7(9): 877–890, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1976:NGC

[BN76] K. S. Banerjee and G. Nagase. A note on the generalisation of Cochran’s theorem. Communications in Statistics: Theory and Methods, 5(9):837–842, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boardman:1977:GCS

[Boa77] Thomas J. Boardman. Graphical contributions to the χ^{2} statistic for two-way contingency tables. Communications in Statistics: Theory and Methods, 6(15):1437–1451, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1978:SUP

[Bof78] Eve Bofinger. Selection using “play-the-winner” sampling for nonnegative observations. Communications in Statistics: Theory and Methods, 7(15):1489–1499, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brunk:1977:LSP

[BP77] H. D. Brunk and Donald A. Pierce. Large sample posterior normality of the population mean. Communications in Statistics: Theory and Methods, 6(1):1–14, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Breslow:1978:PHM

[Bre78] Norman Breslow. The proportional hazards model: applications in epidemiology. Communications in Statistics: Theory and Methods, 7(4):315–332, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1977:FFV

[Bro77a] Lyle D. Broemeling. Forecasting future values of changing sequences. Communications in Statistics: Theory and Methods, 6(1): 87–102, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1977:EDC

[Bro77b] K. G. Brown. On estimation of diagonal covariance matrices by MINQUE. Communications in Statistics: Theory and Methods, 6 (5):471–484, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1978:SIV

[Bro78a] L. D. Broemeling. Simultaneous inferences for variance ratios of some mixed linear models. Communications in Statistics: Theory and Methods, 7(3):297–305, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1978:ESV

[Bro78b] G. H. Brown. Estimation from smallest values. Communications in Statistics: Theory and Methods, 7(1):57–64, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1978:RER

[Bro78c] K. G. Brown. On ridge estimation in rank deficient models. Communications in Statistics: Theory and Methods, 7(2):187–192, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brunden:1978:PCG

[Bru78] Marshall N. Brunden. The probability contours and a goodness-of-fit test for the singly truncated bivariate normal distribution. Communications in Statistics: Theory and Methods, 7(6):557–572, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brucker:1979:NBN

[Bru79] J. Brucker. A note on the bivariate normal distribution. Communications in Statistics: Theory and Methods, 8(2):175–177, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhapkar:1976:MCM

[BS76] V. P. Bhapkar and Grant W. Somes. Multiple comparisons of matched proportions. Communications in Statistics: Theory and Methods, 5(1):17–25, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brogan:1978:EOM

[BS78] Donna R. Brogan and Joseph Sedransk. Estimating one mean of a bivariate normal distribution using a preliminary test of significance and a two-stage sampling scheme. Communications in Statistics: Theory and Methods, 7(2):173–186, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bradley:1979:SAS

[BSL79] Ralph A. Bradley, Sushil S. Srivastava, and Adolf Lanzdorf. Some approaches to statistical analysis of a weather modification experiment. Communications in Statistics: Theory and Methods, 8(11): 1049–1081, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barton:1979:SCM

[BT79] Russell R. Barton and Bruce W. Turnbull. A survey of covariance models for censored life data with an application to recidivism analysis. Communications in Statistics: Theory and Methods, 8(8):723– 750, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhapkar:1979:GTI

[BV79] V. P. Bhapkar and T. C. Venable. On Gart’s test for interaction of second order. Communications in Statistics: Theory and Methods, 8(5):463–470, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burr:1978:PES

[BY78] P. C. Burr and D. H. Young. The power of the exponential scores test for an ordinary renewal process against trend alternatives. Communications in Statistics: Theory and Methods, 7(5):461–473, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Camber:1979:CRP

[Cam79a] Heather Camber. Confidence regions for periodic functions. Communications in Statistics: Theory and Methods, 8(14):1437–1446, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Camber:1979:COS

[Cam79b] Heather A. Camber. Choice of an optimal shape parameter when smoothing noisy data. Communications in Statistics: Theory and Methods, 8(14):1425–1435, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Colvert:1976:EPW

[CB76] Robert E. Colvert and Thomas J. Boardman. Estimation in the piece-wise constant hazard rate model. Communications in Statistics: Theory and Methods, 5(11):1013–1029, 1976. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cima:1976:OCS

[CH76] Joseph A. Cima and Yosef Hochberg. On optimality criteria in simultaneous interval estimation. Communications in Statistics: Theory and Methods, 5(9):875–882, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chanda:1976:SCS

[Cha76] Kamal C. Chanda. Some comments on sample quantiles for dependent observations. Communications in Statistics: Theory and Methods, 5(14):1385–1392, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1978:NMO

[Che78] Ching-Shui Cheng. A note on (m.s)-optimality. Communications in Statistics: Theory and Methods, 7(14):1327–1338, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chechile:1979:EAD

[Che79] R. A. Chechile. Exact and approximate distributions for linear contrasts of means and variances. Communications in Statistics: Theory and Methods, 8(5):447–461, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chosh:1978:CBD

[Cho78] Pankaj Chosh. A characterization of a bimodal distribution. Communications in Statistics: Theory and Methods, 7(5):475–477, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Conover:1976:SAP

[CI76] W. J. Conover and R. L. Iman. On some alternative procedures using ranks for the analysis of experimental designs. Communications in Statistics: Theory and Methods, 5(14):1349–1368, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Conover:1976:CAE

[CK76] W. J. Conover and K. E. Kemp. Comparisons of the asymptotic efficiencies of two sample tests for discrete distributions. Communications in Statistics: Theory and Methods, 5(1):1–15, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chanda:1978:SNE

[CK78] Kamal C. Chanda and Richard W. Kulp. On some nonparametric estimators for the linear Markov scheme. Communications in Statistics: Theory and Methods, 7(5):427–439, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clarke:1979:HNN

[CLB79] William R. Clarke, Peter A. Lachenbruch, and Barabara Broffitt. How non-normality affects the quadratic discriminant function. Communications in Statistics: Theory and Methods, 8(13):1285– 1301, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clickner:1976:LSE

[Cli76] Robert P. Clickner. On least squares estimation for categorical data. Communications in Statistics: Theory and Methods, 5(11):1059– 1064, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clave:1977:PHC

[CM77] J. T. Mc Clave and R. G. Marks. Predicting hospital census using time series regression methods. Communications in Statistics: Theory and Methods, 6(12):1187–1205, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cohen:1977:GSE

[Coh77] Claude Cohen. The generalized symmetric eigenproblem in multivariate statistical models. Communications in Statistics: Theory and Methods, 6(3):277–288, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crow:1977:MVU

[Cro77] Edwin L. Crow. Minimum variance unbiased estimators of the ratio of means of two lognormal variates. Communications in Statistics: Theory and Methods, 6(10):967–975, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaffin:1977:POA

[CSH77] Wilkie W. Chaffin, John T. Sennetti, and Klaus Hinkelmann. A procedure for optimal augmentation of multidimensional designs. Communications in Statistics: Theory and Methods, 6(7):679–689, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cune:1977:CFE

[Cun77] E. D. Mc Cune. On Cornish–Fisher expansions as a ratio of determinants without cumulants. Communications in Statistics: Theory and Methods, 6(3):243–250, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahiya:1977:ETB

[Dah77] Ram C. Dahiya. Estimation in a truncated bivariate Poisson distribution. Communications in Statistics: Theory and Methods, 6 (2):113–120, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1978:TSW

[Dav78] C. E. Davis. A two sample Wilcoxon test for progressively censored data. Communications in Statistics: Theory and Methods, 7(4):389– 398, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dayal:1977:BPT

[Day77] Hari H. Dayal. Bayesian procedures for testing partial equality of two populations when parameters are exchangeable. Communications in Statistics: Theory and Methods, 6(14):1393–1397, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Devore:1977:NRR

[Dev77] Jay L. Devore. A note on the randomized response technique. Communications in Statistics: Theory and Methods, 6(15):1525–1529, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1977:SWD

[DG77] A. Dey and S. C. Gupta. Singular weighing designs and estimation of total weight. Communications in Statistics: Theory and Methods, 6(3):289–295, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1978:BAR

[DG78] Norman R. Draper and Irwin Guttman. Bayesian analysis of reliability in multicomponent stress-strength models. Communications in Statistics: Theory and Methods, 7(5):441–451, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutter:1979:ELM

[DG79] R. Dutter and I. Guttman. On estimation in the linear model when spurious observations are present: a Bayesian approach. Communications in Statistics: Theory and Methods, 8(7):611–635, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1979:ARL

[DGL79] Norman B. Draper, Irwin Guttman, and Lillian Lapczak. Actual rejection levels in a certain stepwise test. Communications in Statistics: Theory and Methods, 8(2):99–105, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donner:1978:SAL

[DH78] Allan Donner and Walter W. Hauck. A simple alternative to the likelihood ratio test in a logistic discrimination problem. Communications in Statistics: Theory and Methods, 7(15):1451–1463, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DiPillo:1976:ABD

[Di 76] Pasquale J. Di Pillo. The application of bias to discriminant analysis. Communications in Statistics: Theory and Methods, 5(9):843– 854, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DiPillo:1977:FAB

[Di 77] Pasquale J. Di Pillo. Further applications of bias to discriminant analysis. Communications in Statistics: Theory and Methods, 6 (10):933–943, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dipillo:1979:BDA

[Dip79] Pasquale J. Dipillo. Biased discriminant analysis: Evaluation of the optimum probability of misclassification. Communications in Statistics: Theory and Methods, 8(14):1447–1457, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dyer:1979:FLC

[DK79] Danny D. Dyer and Jerome P. Keating. A further look at the comparison of normal percentile estimators. Communications in Statistics: Theory and Methods, 8(1):1–16, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dyer:1979:RBE

[DKH79] Dany Dyer, Jerome P. Keating, and Onas L. Hensley. On the relative behavior of estimators of reliability/survivability. Communications in Statistics: Theory and Methods, 8(5):399–416, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Denby:1977:RRE

[DL77] L. Denby and W. A. Larsen. Robust regression estimators compared via Monte Carlo. Communications in Statistics: Theory and Methods, 6(4):335–362, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Decell:1976:FCB

[DM76] Henry P. Decell, Jr. and Salma K. Marani. Feature combinations and the Bhattacharyya criterion. Communications in Statistics: Theory and Methods, 5(12):1143–1152, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dietrich:1979:AFQ

[DM79] F. H. Dietrich and R. G. Marks. Analysis of a factorial quantal response assay using inverse regression. Communications in Statistics: Theory and Methods, 8(1):85–98, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Desu:1977:TSP

[DNV77] M. M. Desu, Subhash C. Narula, and Bernardo Villarreal. A twostage procedure for selecting the best of k exponential distributions. Communications in Statistics: Theory and Methods, 6(12):1223– 1230, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donald:1977:SSP

[Don77] Gary C. Mc Donald. Subset selection procedures based on sample ranges. Communications in Statistics: Theory and Methods, 6(11): 1055–1079, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Downton:1976:SES

[Dow76] F. Downton. Statistics in English schools for the 16–18 + age group. Communications in Statistics: Theory and Methods, 5(10):901–924, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1978:USS

[DQ78] C. E. Davis and Dana Quade. U-statistics for skewness or symmetry. Communications in Statistics: Theory and Methods, 7(5):413–418, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drane:1976:TRR

[Dra76] Wanzer Drane. N the theory of randomized responses to two sensitive questions. Communications in Statistics: Theory and Methods, 5(6):565–574, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Decell:1976:OLC

[DS76a] Henry P. Decell, Jr. and W. G. Smiley III. Optimal linear combinations using Householder transformations. Communications in Statistics: Theory and Methods, 5(12):1153–1162, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutt:1976:IRT

[DS76b] J. E. Dutt and A. P. Soms. An integral representation technique for calculating general multivariate probabilities with an application to multivariate χ^{2}. Communications in Statistics: Theory and Methods, 5(4):377–388, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dodge:1977:EPL

[DS77a] Yadolah Dodge and K. R. Shah. Estimation of parameters in latin squares and graeco-latin squares with missing observations. Communications in Statistics: Theory and Methods, 6(15):1465–1472, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Doviak:1977:UAV

[DS77b] M. J. Doviak and R. L. Scheaffer. Utilizing auxiliary variables in sequential ratio and regression estimation schemes. Communications in Statistics: Theory and Methods, 6(5):391–404, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dwivedi:1978:MMS

[DS78] T. D. Dwivedi and V. K. Srivastava. On the minimum mean squared error estimators in a regression model. Communications in Statistics: Theory and Methods, 7(5):487–494, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duran:1976:SNT

[Dur76] Benjamin S. Duran. A survey of nonparametric tests for scale. Communications in Statistics: Theory and Methods, 5(14):1287–1312, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeWet:1979:ERH

[DV79] T. De Wet and J. W. J. Van Wyk. Efficiency and robustness of Hogg’s adaptive trimmed means. Communications in Statistics:

Theory and Methods, 8(2):117–128, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Helbawy:1977:TCP

[EHB77] Abdalla T. El-Helbawy and Ralph A. Bradley. Treatment contrasts in paired comparisons: convergence of a basic iterative scheme for estimation. Communications in Statistics: Theory and Methods, 6 (3):197–207, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elliott:1979:PMC

[Ell79] Robert D. Elliott. Physically meaningful covariates physically meaningful covariates. Communications in Statistics: Theory and Methods, 8(11):1111–1127, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farebrother:1978:ERC

[Far78] R. W. Farebrother. Estimating regression coefficients under conditional specification: comment. Communications in Statistics: Theory and Methods, 7(2):193–196, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fedorov:1978:SEP

[Fed78] V. V. Fedorov. Some extremal problems in designing discriminating experiments. Communications in Statistics: Theory and Methods, 7 (14):1339–1345, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Federer:1979:PSI

[Fed79] Walter T. Federer. Preface to special issue on covariance. Communications in Statistics: Theory and Methods, 8(8):719–721, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fraser:1976:SCD

[FGS76] D. A. S. Fraser, Irwin Guttman, and George P. H. Styan. Serial correlation and distributions on the sphere. Communications in Statistics: Theory and Methods, 5(2):97–118, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frane:1976:FAT

[FH76] James W. Frane and Maryann Hill. Factor analysis as a tool for data analysis. Communications in Statistics: Theory and Methods, 5(6):487–506, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fomby:1979:MVP

[FH79] Thomas E. Fomby and R. Carter Hill. Multicollinearity and the value of a priori information. Communications in Statistics: Theory and Methods, 8(5):477–486, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fligner:1976:SDF

[FHK76] Michael A. Fligner, Robert V. Hogg, and Timothy J. Killeen. Some distribution-free Rane-like statistics having the Mann–Whitney– Wilcoxon null distribution. Communications in Statistics: Theory and Methods, 5(4):373–376, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fieger:1977:CMN

[Fie77a] W. Fieger. A characterization of the multivariate normal distribution. Communications in Statistics: Theory and Methods, 6(2): 135–140, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Field:1977:CSP

[Fie77b] C. A. Field. Confidence sets as programming problems. Communications in Statistics: Theory and Methods, 6(4):381–390, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fomby:1977:MER

[FJ77] Thomas B. Fomby and S. R. Johnson. MSE evaluation of ridge estimators based on stochastic prior information. Communications in Statistics: Theory and Methods, 6(13):1245–1258, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fochberr:1978:PSI

[FM78] Yosef Fochberr and Puth Marcus. On partitioning successive increments in means of ratios of variances in a chain of normal populations. Communications in Statistics: Theory and Methods, 7 (15):1501–1513, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fligner:1977:CTR

[FPS77] Michael A. Fligner, George E. Policello II, and Jagbir Singh. A comparison of two randomized response survey methods with consideration for the level of respondent protection. Communications in Statistics: Theory and Methods, 6(15):1511–1524, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fraser:1976:TIE

[FR76] W. Fraser and B. L. Raktoe. A theorem on invariance of estimability of linear parametric functions in linear models. Communications in Statistics: Theory and Methods, 5(11):977–983, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gabriel:1979:SSI

[Gab79] K. Ruben Gabriel. Some statistical issues in weather experimentation. Communications in Statistics: Theory and Methods, 8(10): 975–1015, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gani:1978:ASB

[Gan78] J. Gani. The analysis of some biological data. Communications in Statistics: Theory and Methods, 7(10):905–915, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gart:1978:ARC

[Gar78] John J. Gart. The analysis of ratios and cross-product ratios of Poisson variates with application to incidence rates. Communications in Statistics: Theory and Methods, 7(10):917–937, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giesbrecht:1978:EVC

[GB78] Francis G. Giesbrecht and Peter M. Burrows. Estimating variance components in hierarchical structures using MINQUE and restricted maximum livelihood. Communications in Statistics: Theory and Methods, 7(9):891–904, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gbur:1979:AVP

[Gbu79] Edward E. Gbur, Jr. Analysis of variance with Poisson responses. Communications in Statistics: Theory and Methods, 8(5):433–445, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1979:GWD

[GC79] A. K. Gupta and A. K. Chattopadhyay. Gammaization and Wishartness of dependent quadratic forms. Communications in Statistics: Theory and Methods, 8(9):945–951, 1979. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guttman:1976:PIO

[GD76] Irwin Guttman and Rudolf Dutter. Procedures for investigating outliers when estimating in the general univariate linear situationnonfull rank case. Communications in Statistics: Theory and Methods, 5(9):819–835, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goldstein:1977:SDV

[GD77] Matthew Goldstein and William R. Dillon. A stepwise discrete variable selection procedure. Communications in Statistics: Theory and Methods, 6(14):1423–1436, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gentleman:1976:IGT

[Gen76] Jane F. Gentleman. Interactive graphics in a terminal-equipped classroom. Communications in Statistics: Theory and Methods, 5 (10):949–967, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindarajulu:1977:EFV

[GG77] Z. Govindarajulu and A. K. Gupta. Estimators for “the fraction of variance explained” in the one-way classification. Communications in Statistics: Theory and Methods, 6(15):1473–1484, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1977:SMD

[GH77] Shanti S. Gupta and Deng-Yuan Huang. Some multiple decision problems in analysis of variance. Communications in Statistics: Theory and Methods, 6(11):1035–1054, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunst:1978:CS

[GH78] Richard F. Gunst and Tsushung A. Hua. Communications in Statistics: 1972–1976. Communications in Statistics: Theory and Methods, 7(3):197–209, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gokhale:1978:MDI

[GK78] D. V. Gokhale and S. Kullback. The minimum discrimination information approach in analyzing categorical data. Communications in Statistics: Theory and Methods, 7(10):987–1005, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gentle:1977:LAV

[GKS77] J. E. Gentle, W. J. Kennedy, and V. A. Sposito. On least absolute values estimation. Communications in Statistics: Theory and Methods, 6(9):839–845, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunst:1976:GMS

[GM76] R. F. Gunst and R. L. Mason. Generalized mean squared error properties of regression estimators. Communications in Statistics: Theory and Methods, 5(15):1501–1508, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gleeson:1978:IRA

[GM78] A. C. Gleeson and C. A. McGilchrist. An introduction to relationship algebras. Communications in Statistics: Theory and Methods, 7(11):1053–1078, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1979:SPE

[GM79] Malay Ghosh and Nitis Mukhopadhyay. Sequential point estimation of the mean when the distribution is unspecified. Communications in Statistics: Theory and Methods, 8(7):637–652, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gordon:1978:MPT

[Gor78] Robert D. Gordon. On monotonicity of power of two-sample rank tests when testing the alternative g is stochastically larger than f. Communications in Statistics: Theory and Methods, 7(6):535– 541, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindarajulu:1976:BSN

[Gov76] Z. Govindarajulu. A brief survey of nonparametric statistics. Communications in Statistics: Theory and Methods, 5(5):429–453, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grain:1976:MDE

[Gra76] Bradford R. Grain. Matrix density estimation. Communications in Statistics: Theory and Methods, 5(1):89–96, 1976. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gardiner:1978:ANC

[GS78] Jodrph C. Gardiner and Pranab Kumar Sen. Asymptotic normality of a class of time-sequential statistics and applications. Communications in Statistics: Theory and Methods, 7(4):373–388, 1978. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gelfand:1976:DBB

[GT76a] Alan E. Gelfand and D. L. Thomas. Discrimination between the binomial and hypergeometric models. Communications in Statistics: Theory and Methods, 5(3):225–240, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1976:RRM

[GT76b] A. K. Gupta and D. S. Tracy. Recurrence relations for the moments of truncated multinormal distribution. Communications in Statistics: Theory and Methods, 5(9):855–865, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunst:1976:RSC

[Gun76] Richard F. Gunst. Reassessing service courses in statistics. Communications in Statistics: Theory and Methods, 5(10):925–934, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1976:DSI

[Gup76] Ramesh C. Gupta. Distribution of the sum of independent generalized logarithmic series variables. Communications in Statistics: Theory and Methods, 5(1):45–48, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1977:MVU

[Gup77a] Ramesh C. Gupta. Minimum variance unbiased estimation in a modified power series distribution and some of its applications. Communications in Statistics: Theory and Methods, 6(10):977–991, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1977:SRP

[Gup77b] Shanti S. Gupta. Selection and ranking procedures: a brief introduction. Communications in Statistics: Theory and Methods, 6(11): 993–1001, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1979:CSD

[Gup79a] Ramesh C. Gupta. On the characterization of survival distributions in reliability by properties of their renewal densities. Communications in Statistics: Theory and Methods, 8(7):685–697, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1979:SCC

[Gup79b] Ramesh C. Gupta. Some counterexamples in the competing risk analysis. Communications in Statistics: Theory and Methods, 8 (15):1535–1540, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1979:WTP

[Gup79c] Ramesh C. Gupta. Waiting time paradox and size biased sampling. Communications in Statistics: Theory and Methods, 8(6):601–607, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guseman:1977:ALF

[GW77] L. F. Guseman, Jr. and Jay R. Walton. An application of linear feature selection to estimation of proportions. Communications in Statistics: Theory and Methods, 6(7):611–617, 1977. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guseman:1978:MEP

[GW78] L. F. Guseman, Jr. and Jay R. Walton. Methods for estimating proportions of convex combinations of normals using linear feature selection. Communications in Statistics: Theory and Methods, 7 (15):1439–1450, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goldstein:1976:TIC

[GWD76] Matthew Goldstein, Edward Wolf, and William Dillon. On a test of independence for contingency tables. Communications in Statistics: Theory and Methods, 5(2):159–169, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herzberg:1978:RCB

[HA78] Agnes M. Herzberg and David F. Andrews. The robustness of chain block designs and coat-of-mail designs. Communications in Statistics: Theory and Methods, 7(5):479–485, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haack:1976:NSC

[Haa76] D. G. Haack. A nonsymbolic statistics course. Communications in Statistics: Theory and Methods, 5(10):943–947, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haack:1978:AAC

[Haa78] Dennis G. Haack. An adaptive approach to compound multiple decisions. Communications in Statistics: Theory and Methods, 7 (3):229–239, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:1978:PLS

[Haq78] M. Safiul Haq. On prediction from the location-scale model with equi correlated response. Communications in Statistics: Theory and Methods, 7(9):857–866, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haralick:1976:TLR

[Har76] Robert M. Haralick. The table look-up rule. Communications in Statistics: Theory and Methods, 5(12):1163–1191, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harter:1977:NLA

[Har77] H. Leon Harter. Nonuniqueness of least absolute values regression. Communications in Statistics: Theory and Methods, 6(9):829–838, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harvey:1978:URR

[Har78] A. C. Harvey. On the unbiasedness of robust regression estimators. Communications in Statistics: Theory and Methods, 7(8):779–783, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoq:1978:DSR

[HAT78] A. K. M. Sirajul Hoq, Mir M. Ali, and J. G. C. Templeton. The distribution of Student’s ratio for samples from an exponential population. Communications in Statistics: Theory and Methods, 7(9): 837–850, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkins:1977:CNS

[Haw77] Douglas M. Hawkins. Comment on “A new statistic for testing suspected outliers”. Communications in Statistics: Theory and Methods, 6(5):435–438, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Tik75, Tik77].

Holbert:1977:BIR

[HB77] Donald Holbert and Lyle Broemeling. Bayesian inferences related to shifting sequences and two-phase regression. Communications in Statistics: Theory and Methods, 6(3):265–275, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:1978:ERC

[HB78] Chien-Pai Han and T. A. Bancroft. Estimating regression coefficients tinder conditional specification. Communications in Statistics: Theory and Methods, 7(1):47–56, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanson:1979:BWM

[HBB^{+}79] Morgan A. Hanson, Lawrence E. Barker, Charles L. Bach, Ldward A. Cooley, and Charles H. Hunter. A bibliography of weather modification experiments. Communications in Statistics: Theory and Methods, 8(11):1129–1153, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henry:1976:NEW

[Hen76] C. E. Mc Henry. A note on expressing Wilks’ A as the product of independent beta variables. Communications in Statistics: Theory and Methods, 5(11):1047–1053, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holt:1976:RPF

[HF76] D. Holt and J. G. Fryer. On the robustness of the power function of the one-sample test for the negative exponential distribution. Communications in Statistics: Theory and Methods, 5(8):723–734, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hill:1977:CSN

[HFJ77] R. Carter Hill, Thomas B. Fomby, and S. R. Johnson. Component selection norms for principal components regression. Communications in Statistics: Theory and Methods, 6(4):309–334, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harkness:1976:ITP

[HG76] William L. Harkness and Ashok V. Godambe. Inequalities for tail probabilities for the multivariate normal distribution. Communications in Statistics: Theory and Methods, 5(7):689–692, 1976. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henderson:1979:ACM

[HH79] Charles R. Henderson, Jr. and C. R. Henderson. Analysis of covariance in mixed models with unequal subclass numbers. Communications in Statistics: Theory and Methods, 8(8):751–787, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hidiroglou:1978:AIM

[Hid78] M. A. Hidiroglou. An approximation of the inverse moments of the positive hypergeometric distribution. Communications in Statistics: Theory and Methods, 7(15):1475–1487, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hill:1979:ECM

[Hil79] Richard W. Hill. On estimating the covariance matrix of robust regression M-estimates. Communications in Statistics: Theory and Methods, 8(12):1183–1196, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoerl:1976:RRI

[HK76a] Arthur E. Hoerl and Robert W. Kennard. Ridge regression iterative estimation of the biasing parameter. Communications in Statistics: Theory and Methods, 5(1):77–88, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hollander:1976:NEB

[HK76b] Myles Hollander and Ramesh M. Korwar. Nonparametric empirical Bayes estimation of the probability that x ≥ y. Communications in Statistics: Theory and Methods, 5(14):1369–1383, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hochberg:1976:TSM

[HL76] Yosef Hochberg and Peter A. Lachenbruch. Two stage multiple comparison procedures based on the Studentized range. Communications in Statistics: Theory and Methods, 5(15):1447–1453, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haack:1976:GUM

[HM76] D. C. Haack and H. E. McKean. Graduate and undergraduate modular instruction. Communications in Statistics: Theory and Methods, 5(10):935–941, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hocking:1979:EID

[HM79] R. R. Hocking and D. L. Marx. Estimation with incomplete data: An improved computational method and the analysis of nested data. Communications in Statistics: Theory and Methods, 8(12):1155– 1181, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harter:1979:ARE

[HMC79] H. Leon Harter, Albert H. Moore, and Thomas F. Curry. Adaptive robust estimation of location and scale parameters of symmetric populations. Communications in Statistics: Theory and Methods, 8 (15):1473–1491, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hensler:1977:GFT

[HMM77] G. L. Hensler, K. G. Mehrotra, and J. E. Michalek. A goodness of fit test for multivariate normality. Communications in Statistics: Theory and Methods, 6(1):33–41, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herrin:1978:UAP

[HN78] G. D. Herrin and J. B. Neuhardt. A useful approximation for planning block designs with potential missing data. Communications in Statistics: Theory and Methods, 7(8):755–766, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:1979:MLI

[HN79] M. Safiul Haq and V. Ming Ng. On marginal likelihood inference for the intra-class correlation coefficient. Communications in Statistics: Theory and Methods, 8(2):179–189, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hochberg:1977:UGP

[Hoc77] Yosef Hochberg. On the use of general positive quadratic forms in simultaneous inference. Communications in Statistics: Theory and Methods, 6(13):1289–1300, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1976:NDN

[Hog76] Robert V. Hogg. A new dimension to nonparametric tests. Communications in Statistics: Theory and Methods, 5(14):1313–1325, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1977:IRP

[Hog77] Robert V. Hogg. An introduction to robust procedures. Communications in Statistics: Theory and Methods, 6(9):789–794, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horn:1976:TSE

[Hor76] Susan Dadakis Horn. Teaching statistics to elementary school students. Communications in Statistics: Theory and Methods, 5(10): 895–899, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1976:MSS

[HP76] Deng-Yuan Huang and S. Panchapakesan. A modified subset selection formulation with special reference to one-way and two-way layout experiments. Communications in Statistics: Theory and Methods, 5(7):621–633, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsuan:1976:GFT

[HR76] A. Hsuan and D. S. Robson. The χ^{2}-goodness-of-fit tests with moment type estimators. Communications in Statistics: Theory and Methods, 5(15):1509–1519, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hughes:1979:SPO

[HR79] Edward Hughes and J. N. K. Rao. Some problems of optimal allocation in sample surveys involving inequality constraints. Communications in Statistics: Theory and Methods, 8(15):1551–1574, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hedayat:1977:EAR

[HRT77] A. Hedayat, B. L. Raktoe, and Prem P. Talwar. Examination and analysis of residuals: A test for detecting a monotonic relation between mean and variance in regression through the origin. Communications in Statistics: Theory and Methods, 6(6):497–506, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsieh:1979:MES

[Hsi79] H. K. Hsieh. On multivariate extension of Sukhatme–Perng–Littell test. Communications in Statistics: Theory and Methods, 8(5):471– 475, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1978:TTS

[HT78] Tea-Yuan Hwang and Jain-Feng Tsay. The treatment of ties in some linear rank tests when the underlying distribution is discrete. Communications in Statistics: Theory and Methods, 7(9):813–827, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hettmansperger:1977:RPS

[HU77] Thomas P. Hettmansperger and Jessica M. Utts. Robustness properties for a simple class of rank estimates. Communications in Statistics: Theory and Methods, 6(9):855–868, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1976:SSP

[Hua76] Deng-Yuan Huang. On some selection procedures for normal means problems with unequal sample sizes. Communications in Statistics: Theory and Methods, 5(15):1489–1499, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huitema:1979:ACM

[Hui79] Bradley E. Huitema. An analysis of covariance Markov model for interrupted time series experiments. Communications in Statistics: Theory and Methods, 8(8):789–797, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hutchinson:1979:VCS

[Hut79] T. P. Hutchinson. The validity of the chi-squared test when expected frequencies are small: a list of recent research references. Communications in Statistics: Theory and Methods, 8(4):327–335, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holland:1977:RRU

[HW77] Paul W. Holland and Roy E. Welsch. Robust regression using iteratively reweighted least-squares. Communications in Statistics:

Theory and Methods, 6(9):813–827, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1976:RDM

[Hwa76] F. K. Hwang. The rank distribution of the maximin in a random matrix. Communications in Statistics: Theory and Methods, 5(15): 1533–1538, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1978:CEW

[Hwa78] Tea-Yuan Hwang. The Chernoff efficiency of the Wilcoxon rank test. Communications in Statistics: Theory and Methods, 7(6):543– 555, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iman:1976:NAE

[ID76] Ronald L. Iman and James M. Davenport. New approximations to the exact distribution of the Kruskal–Wallis test statistic. Communications in Statistics: Theory and Methods, 5(14):1335–1348, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iman:1976:AED

[Ima76] Ronald L. Iman. An approximation to the exact distribution of the Wilcoxon–Mann–Whitney rank sum test statistic. Communications in Statistics: Theory and Methods, 5(7):587–598, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Imrey:1979:ACD

[ISF79] Peter B. Imrey, E. Sobel, and M. E. Francis. Analysis of categorical data obtained by stratified random sampling. Communications in Statistics: Theory and Methods, 8(7):653–670, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

James:1979:GIF

[Jam79] M. James. The generalised inverse form of canonical correlation. Communications in Statistics: Theory and Methods, 8(6):561–568, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1976:RET

[JC76] Douglas H. Jones and Lewis Coopersmith. A ratio estimator of the total of a subpopulation. Communications in Statistics: Theory and Methods, 5(3):251–260, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1977:SGF

[JK77] N. L. Johnson and S. Kotz. On some generalized Farlie–Gumbel– Morgenstern distributions — II regression, correlation and further generalizations. Communications in Statistics: Theory and Methods, 6(6):485–496, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1977:SSG

[Joh77] Peter W. M. John. Series of semiregular group divisible designs. Communications in Statistics: Theory and Methods, 6(14):1385– 1392, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1978:SUD

[Joh78] Thomes Johnson. Selection without (unfair) discrimination. Communications in Statistics: Theory and Methods, 7(11):1079–1098, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1977:OSA

[Jon77a] Douglas H. Jones. A one-sample adaptive distribution-free test with a stable power function. Communications in Statistics: Theory and Methods, 6(9):869–877, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1977:FSR

[Jon77b] E. R. Jones. Fixed-size rectangular confidence regions. Communications in Statistics: Theory and Methods, 6(3):251–264, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johari:1976:PD

[JS76] Shyaai Johari and Stanley L. Selove. Partitioning a distribution. Communications in Statistics: Theory and Methods, 5(2):133–147, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1978:GMT

[JS78] Suneeta Joshi and Y. S. Sathe. A generalization of Mosteller’s test. Communications in Statistics: Theory and Methods, 7(7):709–715, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabir:1979:ELS

[KA79] A. B. M. Lutful Kabir and M. Ahsanullah. Estimation of the location and scale parameters of a power function distribution based on a selected number of order statistics. Communications in Statistics: Theory and Methods, 8(2):139–149, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadane:1976:WUT

[Kad76] Joseph B. Kadane. For what use are tests of hypotheses and tests of significance, introduction. Communications in Statistics: Theory and Methods, 5(8):735–736, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadiyala:1979:ORR

[Kad79] Krishna Kadiyala. Operational ridge regression estimators under the prediction goal. Communications in Statistics: Theory and Methods, 8(14):1377–1391, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kambo:1978:MLE

[Kam78] N. S. Kambo. Maximum likelihood estimators of the location and scale parameters of the exponential distribution from a censored sample. Communications in Statistics: Theory and Methods, 7(12): 1129–1132, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khuri:1977:SDC

[KC77] A. I. Khuri and J. A. Cornell. Secondary design considerations for minimum bias estimation. Communications in Statistics: Theory and Methods, 6(7):631–647, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1979:NMM

[KC79] Ashish Kumar and P. C. Consul. Negative moments of a modified power series distribution and bias of the maximum likelihood estimator. Communications in Statistics: Theory and Methods, 8(2): 151–166, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kedem:1977:ANL

[Ked77] B. Kedem. On the asymptotic normality of linear rank statistics. Communications in Statistics: Theory and Methods, 6(3):221–229, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koch:1976:ACS

[KEKL76] Gary G. Koch, M. Nabil El-Khorazaty, and Alcinda L. Lewis. The asymptotic covariance structure of log-linear model estimated parameters for the multiple recapture census. Communications in Statistics: Theory and Methods, 5(14):1425–1445, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kempthore:1976:WUT

[Kem76] Oscar Kempthore. Of what use are tests of significance and tests of hypothesis. Communications in Statistics: Theory and Methods, 5(8):763–777, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:1978:SSP

[Kem78] Adrienne W. Kemp. Slustep size probabilities for generalized Poisson distributions. Communications in Statistics: Theory and Methods, 7(15):1433–1438, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanarek:1978:GFC

[KF78] Paula Kanarek and Lloyd Fisher. A goodness-of-fit comparison of four survival curve models. Communications in Statistics: Theory and Methods, 7(5):419–426, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kullback:1978:IAL

[KG78] S. Kullback and D. V. Gokhale. Information analysis of linear interactions in contingency tables. Communications in Statistics: Theory and Methods, 7(1):27–45, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:1978:NTM

[Kha78] Rasul A. Khan. A note on testing the means of normal distributions with known coefficient of variations. Communications in Statistics: Theory and Methods, 7(9):867–876, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khuri:1978:CSS

[Khu78] A. I. Khuri. A conservative sample size for the comparison of several proportions. Communications in Statistics: Theory and Methods, 7 (13):1283–1293, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kiefer:1978:AAF

[Kie78] J. Kiefer. Asymptotic approach to families of design problems. Communications in Statistics: Theory and Methods, 7(14):1347–1362, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kochar:1978:CDF

[Koc78] Subhash C. Kochar. A class of distribution-free tests for the twosample slippage problem. Communications in Statistics: Theory and Methods, 7(13):1243–1252, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koopmans:1978:ACP

[Koo78] L. H. Koopmans. An application of a crossing probability bound to population biology. Communications in Statistics: Theory and Methods, 7(2):127–137, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korwar:1977:CLP

[Kor77] R. M. Korwar. On characterizing Lagrangian, Poisson, and quasibinomial distributions. Communications in Statistics: Theory and Methods, 6(14):1409–1415, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koul:1977:TNB

[Kou77] Hira L. Koul. A test for new better than used. Communications in Statistics: Theory and Methods, 6(6):563–573, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koul:1978:TNB

[Kou78] Hira L. Koul. Testing for new is better than used in expectation. Communications in Statistics: Theory and Methods, 7(7):685–701, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koziol:1979:TBS

[Koz79] James A. Koziol. A test for bivariate symmetry based on the empirical distribution function. Communications in Statistics: Theory and Methods, 8(3):207–221, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1978:STL

[KP78] K. Ravi Kumar and Elinor S. Pape. Selecting a test level for random effects. Communications in Statistics: Theory and Methods, 7(7): 703–707, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koziol:1977:MCA

[KR77] James A. Koziol and Nancy Reid. On multiple comparisons among K samples subject to unequal patterns of censorship. Communications in Statistics: Theory and Methods, 6(12):1149–1164, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatri:1979:EMP

[KR79] G. G. Khatri and R. T. Ratani. On estimation of the mean parameter of a truncated normal distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 8(3):237–244, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krewski:1978:JSF

[Kre78] Daniel Krewski. Jackknifing U-statistics in finite populations. Communications in Statistics: Theory and Methods, 7(1):1–12, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnan:1976:NBC

[Kri76] Marakatha Krishnan. The noncentral bivariate chi-squared distribution and extensions. Communications in Statistics: Theory and Methods, 5(7):647–660, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kudo:1978:TED

[KT78] Alio Kudô and Tomoyuki Tarumi. 2 × 2 tables emerging out of different chance mechanisms. Communications in Statistics: Theory and Methods, 7(10):977–986, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kurotschka:1978:ODC

[Kur78] V. Kurotschka. Optimal design of complex experiments with qualitative factors of influence. Communications in Statistics: Theory and Methods, 7(14):1363–1378, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khursheed:1977:SBP

[KW77] Alam Khursheed and K. T. Wallenius. On selecting the best population. Communications in Statistics: Theory and Methods, 6(11): 1091–1104, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1976:LSS

[LA76] Austin F. S. Lee and Ralph B. D. Agostino. Levels of significance of some two-sample tests when observations are from compound normal distributions. Communications in Statistics: Theory and Methods, 5(4):325–342, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawless:1978:RRE

[Law78] J. F. Lawless. Ridge and related estimation procedures: theory and practice. Communications in Statistics: Theory and Methods, 7(2):139–164, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemeshow:1977:PBH

[LE77] Stanley Lemeshow and Robert Epp. Properties of the balanced halfsample and jackknife variance estimation techniques in the linear case. Communications in Statistics: Theory and Methods, 6(13): 1259–1274, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ledolter:1979:RAP

[Led79] Johannes Ledolter. A recursive approach to parameter estimation in regression and time series models. Communications in Statistics: Theory and Methods, 8(12):1227–1245, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lenth:1977:RS

[Len77] Russell V. Lenth. Robust splines. Communications in Statistics: Theory and Methods, 6(9):847–854, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leonard:1977:BSE

[Leo77] Tom Leonard. Bayesian simultaneous estimation for several multinomial distributions. Communications in Statistics: Theory and Methods, 6(7):619–630, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lepage:1976:APE

[Lep76] Yves Lepage. Asymptotic power efficiency for a location and scale problem. Communications in Statistics: Theory and Methods, 5 (13):1257–1274, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lepage:1977:CNT

[Lep77] Yves Lepage. A class of nonparametric tests for location and scale parameters. Communications in Statistics: Theory and Methods, 6(7):649–659, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyons:1979:DPS

[LH79] N. I. Lyons and K. Hutcheson. Distributional properties of Simpson’s index of diversity. Communications in Statistics: Theory and Methods, 8(6):569–574, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liebetrau:1977:TRT

[Lie77] Albert M. Liebetrau. Tests of randomness in two dimensions. Communications in Statistics: Theory and Methods, 6(14):1367–1383, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lingappaiah:1977:SIG

[Lin77] G. S. Lingappaiah. On some inflated generalized discrete distributions. Communications in Statistics: Theory and Methods, 6(3): 231–241, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lingappaiah:1979:BAP

[Lin79] G. S. Lingappaiah. Bayesian approach to the prediction problem in complete and censored samples from the gamma and exponential populations. Communications in Statistics: Theory and Methods, 8 (14):1403–1423, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawless:1976:THE

[LM76] Jerald F. Lawless and Nancy R. Mann. Tests for homogeneity of extreme value scale parameters. Communications in Statistics: Theory and Methods, 5(5):389–405, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laubscher:1976:CIS

[LO76] N. F. Laubscher and Robert E. Odeh. A confidence interval for the scale parameter based on Sukhtmes two-sample statistic. Communications in Statistics: Theory and Methods, 5(14):1393–1407, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Locke:1976:TCH

[Loc76] Charles Locke. A test for the composite hypothesis that a population has a gamma distribution. Communications in Statistics: Theory and Methods, 5(4):351–364, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyons:1977:TRF

[LP77] N. I. Lyons and C. H. Proctor. A test for regression function adequacy. Communications in Statistics: Theory and Methods, 6(1): 81–86, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

LaCava:1976:SES

[LR76] Gerald J. La Cava and Kenneth Reid. Scale estimators for symmetric stable distributions. Communications in Statistics: Theory and Methods, 5(15):1521–1531, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Locke:1978:TU

[LS78] Charles Locke and John D. Spurrier. On tests of uniformity. Communications in Statistics: Theory and Methods, 7(3):241–258, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lombard:1979:AST

[LS79] F. Lombard and J. W. H. Swanepoel. An asymptotic sequential test based on confidence sequences. Communications in Statistics: Theory and Methods, 8(2):107–116, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawless:1976:SSR

[LW76] J. F. Lawless and P. Wang. A simulation study of ridge and other regression estimators. Communications in Statistics: Theory and Methods, 5(4):307–323, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lynch:1978:DTA

[Lyn78] G. W. Lynch. A decision theoretic approach to ridits. Communications in Statistics: Theory and Methods, 7(6):607–614, 1978. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Milliken:1977:TDG

[MA77] George A. Milliken and Fikri Akdeniz. A theorem on the difference of the generalized inverses of two nonnegative matrices. Communications in Statistics: Theory and Methods, 6(1):73–79, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mann:1978:CSS

[Man78] Nancy R. Mann. Calculation of small-sample Weibull tolerance bounds for accelerated testing. Communications in Statistics: Theory and Methods, 7(2):97–112, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcus:1978:FRS

[Mar78a] Ruth Marcus. Further results on simultaneous confidence bounds in normal models with restricted alternatives. Communications in Statistics: Theory and Methods, 7(6):573–590, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mardia:1978:SPC

[Mar78b] K. V. Mardia. Some properties of classical multi-dimensional scaling. Communications in Statistics: Theory and Methods, 7(13): 1233–1241, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathai:1979:NND

[Mat79] A. M. Mathai. On the non-null distributions of test statistics connected with exponential populations. Communications in Statistics: Theory and Methods, 8(1):47–55, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mantel:1978:MHS

[MB78] Nathan Mantel and David P. Byar. Marginal homogeneity, symmetry, and independence. Communications in Statistics: Theory and Methods, 7(10):953–976, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielke:1976:MRP

[MBJ76] Paul W. Mielke, Jr., Kenneth J. Berry, and Earl S. Johnson. Multiresponse permutation procedures for a priori classifications. Communications in Statistics: Theory and Methods, 5(14):1409–1424, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McClave:1978:EOM

[McC78] James T. McClave. Estimating the order of moving average models: the max χ^{2} method. Communications in Statistics: Theory and Methods, 7(3):259–276, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1976:ADN

[MCCC76] Govind S. Mudholkar, Yogendra P. Chaubey, and Lin Ching-Chuong. Approximations for the doubly noncentral-F distribution. Communications in Statistics: Theory and Methods, 5(1):49–63, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McLachlan:1977:CSD

[McL77] G. J. McLachlan. Constrained sample discrimination with the Studentized classification statistic w. Communications in Statistics: Theory and Methods, 6(6):575–583, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McLachlan:1979:CEP

[McL79] G. J. McLachlan. A comparison of the estimative and predictive methods of estimating posterior probabilities. Communications in Statistics: Theory and Methods, 8(9):919–929, 1979. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Milliken:1978:PTH

[MD78] George A. Milliken and Robert L. Debruin. A procedure to test hypotheses for nonlinear models. Communications in Statistics: Theory and Methods, 7(1):65–79, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maric:1979:SSC

[MG79] Nebojsa Marić and Franklin A. Grayhi. Small samples confidencce intervals on common mean of two normal distributions variances. Communications in Statistics: Theory and Methods, 8(13):1255– 1269, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McKean:1976:THB

[MH76] Joseph W. McKean and Thomas P. Hettmansperger. Tests of hypotheses based on ranks in the general linear model. Communications in Statistics: Theory and Methods, 5(8):693–709, 1976. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mantel:1978:LRA

[MH78] Nathan Mantel and Benjamin F. Hankey. A logistic regression analysis of response-time data where the hazard function is time dependent. Communications in Statistics: Theory and Methods, 7(4): 333–347, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McAllister:1979:ESL

[MH79] Paul R. McAllister and Burt S. Holland. On O.L.S. estimation in singular linear models with general positive definite dispersion matrix. Communications in Statistics: Theory and Methods, 8(3): 299–302, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Michalek:1976:UCE

[Mic76] Joel Michalek. Uniform convergence of an empirical cdf based on a relatively small number of order statistics. Communications in Statistics: Theory and Methods, 5(3):241–250, 1976. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielke:1979:ANN

[Mie79a] P. W. Mielke. On asymptotic non-normality of null distributions of MRPP statistics. Communications in Statistics: Theory and Methods, 8(15):1541–1550, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielke:1979:SPN

[Mie79b] P. W. Mielke. Some parametric, nonparametric and permutation inference procedures result ind from weather modification experiments. Communications in Statistics: Theory and Methods, 8(11): 1083–1096, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mieseke:1979:BSS

[Mie79c] Klaus-J. Mieseke. Bayesian subset selection for additive and linear loss function. Communications in Statistics: Theory and Methods, 8 (12):1205–1226, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehrotra:1977:LMP

[MJB77] K. G. Mehrotra, R. A. Johnson, and G. K. Bhattacharyya. Locally most powerful rank tests for multiple-censored data. Communications in Statistics: Theory and Methods, 6(5):459–469, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehrotra:1979:EFI

[MJB79] Kishan G. Mehrotra, R. A. Johnson, and G. K. Bhattacharyya. Exact fisher information for censored samples and the extended hazard rate functions. Communications in Statistics: Theory and Methods, 8(15):1493–1510, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meydrech:1976:QAC

[MK76] Edward F. Meydrech and Lawrence L. Kupper. On quadratic approximation of a cubic response. Communications in Statistics: Theory and Methods, 5(13):1205–1213, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Myers:1979:NPS

[MK79] R. H. Myers and A. I. Khuri. A new procedure for steepest ascent. Communications in Statistics: Theory and Methods, 8(14):1359– 1376, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marks:1978:MTG

[MR78] R. G. Marks and P. V. Rao. A modified Tiao–Guttman rule for multiple outliers. Communications in Statistics: Theory and Methods, 7(2):113–126, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McDonald:1979:APE

[MR79] James B. McDonald and Michael R. Ransom. Alternative parameter estimators based upon grouped data. Communications in Statistics: Theory and Methods, 8(9):899–917, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Majumdar:1977:ROT

[MS77] Hiranmay Majumdar and Pranab Kumar Sen. Rank order tests for grouped data under progressive censoring. Communications in Statistics: Theory and Methods, 6(6):507–524, 1977. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Majumdar:1978:IPJ

[MS78a] Hiranmay Majumdar and Pranab Kumar Sen. Invariance principles for jackknifing U-statistics for finite population sampling and some applications. Communications in Statistics: Theory and Methods, 7 (11):1007–1025, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Majumdar:1978:NTS

[MS78b] Hiranmay Majumdar and Pranab Kumar Sen. Nonparametric testing for simple regression under progressive censoring with staggering entry and random withdrawal. Communications in Statistics: Theory and Methods, 7(4):349–371, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1979:ARC

[MSVS79] A. J. Miller, D. E. Shaw, L. G. Veitch, and E. J. Smith. Analyzing the results of a cloud-seeding experiment in Tasmania. Communications in Statistics: Theory and Methods, 8(10):1017–1047, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muirhead:1976:FDE

[Mui76] Robb J. Muirhead. Forms for the distribution of the ellipticity statistic for bivariate sphericity. Communications in Statistics: Theory and Methods, 5(5):413–420, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1979:SCT

[Muk79] Nitis Mukhopadhyay. Some comments on two-stage selection procedures. Communications in Statistics: Theory and Methods, 8(7): 671–683, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:1976:BEV

[Nai76] K. Aiyappan Nair. Bivariate extreme value distributions. Communications in Statistics: Theory and Methods, 5(6):575–581, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naik:1977:SSS

[Nai77] Umesh D. Naik. Some subset selection problems. Communications in Statistics: Theory and Methods, 6(10):955–966, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nakicenovic:1979:ANC

[Nak79] Nebojsa Nakicenovic. Annex: Numerical calculations. Communications in Statistics: Theory and Methods, 8(3):231–235, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagase:1976:DSD

[NB76] G. Nagase and K. S. Banerjee. On the distribution of a second degree polynomial statistic in singular normal variates. Communications in Statistics: Theory and Methods, 5(5):407–412, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagase:1978:IBT

[NB78] G. Nagase and K. S. Banerjee. On the independence between two generalized second degree polynomial statistics and their covariance. Communications in Statistics: Theory and Methods, 7(8):733–738, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelson:1979:LST

[Nel79] Paul I. Nelson. Least squares tests for discrete parameter stochastic processes. Communications in Statistics: Theory and Methods, 8 (3):283–297, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neyman:1976:TSH

[Ney76] Jerzy Neyman. Tests of statistical hypotheses and their use in studies of natural phenomena. Communications in Statistics: Theory and Methods, 5(8):737–751, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neyman:1979:DPM

[Ney79] Jerzy Neyman. Developments in probability and mathematical statistics generated by studies in meteorology and weather modification. Communications in Statistics: Theory and Methods, 8 (11):1097–1110, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nordbrock:1976:SDB

[Nor76] Earl Nordbrock. A sequential design for binomial clinical trials. Communications in Statistics: Theory and Methods, 5(15):1469– 1478, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neff:1976:FRE

[NS76] Norman Neff and William E. Strawderman. Further remarks on estimating the parameter of a noncentral chi-square distribution. Communications in Statistics: Theory and Methods, 5(1):65–76, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odell:1976:CSM

[OB76] P. L. Odell and J. P. Basu. Concerning several methods for estimating crop acreages using remote sensing data. Communications in Statistics: Theory and Methods, 5(12):1091–1114, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OBryan:1979:RCM

[O’B79] T. O’Bryan. Rates of convergence in a modified empirical Bayes estimation problem involving Poisson distributions. Communications in Statistics: Theory and Methods, 8(2):167–174, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Odell:1976:CPR

[Ode76] Patrick L. Odell. Current and past roles of the statistician in space applications. Communications in Statistics: Theory and Methods, 5 (12):1077–1089, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oliker:1978:RBS

[Oli78] V. I. Oliker. On the relationship between the sample size and the number of variables in a linear regression model. Communications in Statistics: Theory and Methods, 7(6):509–516, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oman:1978:BCS

[Oma78] Samuel D. Oman. A Bayesian comparison of some estimators used in linear regression with multicollinear data. Communications in Statistics: Theory and Methods, 7(6):517–534, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Padmanabhan:1977:HLE

[Pad77] A. R. Padmanabhan. Hodges–Lehmann estimators in the case of grouped data — I. Communications in Statistics: Theory and Methods, 6(4):371–380, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1979:SEN

[Pan79] B. N. Pandey. On shrinkage estimation of normal population variance. Communications in Statistics: Theory and Methods, 8(4): 359–365, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1976:PSD

[Par76] C. J. Park. The power series distribution with unknown truncation parameter: Two sample problem. Communications in Statistics: Theory and Methods, 5(15):1539–1544, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peters:1976:NEM

[PC76] Charles Peters and William A. Coberly. The numerical evaluation of the maximum-likelihood estimate of mixture proportions. Communications in Statistics: Theory and Methods, 5(12):1127–1135, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peele:1979:CAP

[Pee79] Lawrence Peele. Characterizations of autoregressive prediction and moving average prediction as discrete-time spline interpolation. Communications in Statistics: Theory and Methods, 8(1):35–46, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perng:1977:AET

[Per77] S. K. Perng. An asymptotically efficient test for the location parameter and the scale parameter of an exponential distribution. Communications in Statistics: Theory and Methods, 6(14):1399–1407, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Petkau:1979:ESL

[Pet79a] A. John Petkau. Exact slopes for a life testing problem involving the two parameter exponential distribution. Communications in Statistics: Theory and Methods, 8(15):1511–1521, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Petkau:1979:LTP

[Pet79b] J. John Petkau. On a life testing problem involving the two parameter exponential distribution. Communications in Statistics: Theory and Methods, 8(15):1523–1534, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pettitt:1979:TBN

[Pet79c] A. N. Pettitt. Testing for bivariate normality using the empirical distribution function. Communications in Statistics: Theory and Methods, 8(7):699–712, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paik:1977:ABN

[PF77] U. B. Paik and W. T. Federer. Analysis of binary number association scheme partially balanced designs. Communications in Statistics: Theory and Methods, 6(10):895–931, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Plachky:1977:NSR

[Pla77] Detlef Plachky. Note on sums of a random number of random variables with exponential family distributions. Communications in Statistics: Theory and Methods, 6(13):1305–1309, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puri:1976:BFE

[PN76] P. D. Puri and A. K. Nigam. Balanced factorial experiments. I. Communications in Statistics: Theory and Methods, 5(7):599–619, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puri:1977:BBD

[PN77a] P. D. Puri and A. K. Nigam. Balanced block designs. Communications in Statistics: Theory and Methods, 6(12):1171–1179, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puri:1977:PEB

[PN77b] P. D. Puri and A. K. Nigam. Partially efficiency balanced designs. Communications in Statistics: Theory and Methods, 6(8):753–771, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puri:1978:BFE

[PN78] P. D. Puri and A. K. Nigam. Balanced factorial experiments. II. Communications in Statistics: Theory and Methods, 7(6):591–605, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pollak:1979:CRE

[Pol79] Moshe Pollak. A class of robust estimators. Communications in Statistics: Theory and Methods, 8(6):509–531, 1979. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Plackett:1978:DMS

[PP78] R. L. Plackett and S. R. Paul. Dirichlet models for square contingency tables. Communications in Statistics: Theory and Methods, 7(10):939–952, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patil:1977:ILR

[PR77] G. P. Patil and M. V. Ratnaparkhi. Invariance of linearity of regression and related characterizations of some classical discrete distributions. Communications in Statistics: Theory and Methods, 6 (2):167–174, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peritz:1978:SSM

[PR78] Eric Peritz and Ray Rose. Some statistical methods for survival analysis with rare failures. Communications in Statistics: Theory and Methods, 7(7):619–636, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pratt:1976:DQW

[Pra76] John W. Pratt. A discussion of the question: for what use are tests of hypotheses and tests of significance. Communications in Statistics: Theory and Methods, 5(8):779–787, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prajneshu:1979:STI

[Pra79] Prajneshu. Stability of two interacting species subject to stochastic parameter variation. Communications in Statistics: Theory and Methods, 8(1):57–69, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Price:1976:NEB

[Pri76] Bertram Price. A note on estimation in Bernoulli trials with dependence. Communications in Statistics: Theory and Methods, 5 (7):661–671, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peters:1976:NOL

[PS76a] B. C. Peters, Jr. and J. L. Solomon. A note on optimum linear feature extraction for Gaussian populations with equal covariances and equal a priori probabilities. Communications in Statistics: Theory and Methods, 5(12):1137–1141, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pinski:1976:CNE

[PS76b] K. Pinski and V. A. Sposito. Consistency of the normal equations. Communications in Statistics: Theory and Methods, 5(11):1055– 1057, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pradhan:1976:MFE

[PS76c] Meena Pradhan and Y. S. Sathe. Modification of fixed even sample size procedure for play-the-winner. Communications in Statistics: Theory and Methods, 5(4):365–371, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Panchapakesan:1977:SSP

[PS77a] S. Panchapakesan and T. J. Santner. Subset selection procedures for Δ_{P} -superior populations. Communications in Statistics: Theory and Methods, 6(11):1081–1090, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papaioannou:1977:RCI

[PS77b] Takis Papaioannou and Ted Speevak. Rank correlation inequalities with missing data. Communications in Statistics: Theory and Methods, 6(1):67–72, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Phadia:1979:EBA

[PS79] E. G. Phadia and V. Susarla. An empirical Bayes approach to twosample problems with censored data. Communications in Statistics: Theory and Methods, 8(13):1327–1351, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Price:1976:PCE

[PT76] Bertram Price and Aaron Tenenbein. Probability computations for an extension of the geometric distribution. Communications in Statistics: Theory and Methods, 5(11):1031–1039, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pukelsheim:1977:ETB

[Puk77] Friedrich Pukelsheim. Equality of two blues and ridge-type estimates. Communications in Statistics: Theory and Methods, 6(7): 603–610, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Padgett:1977:BER

[PW77] W. J. Padgett and L. J. Wei. Bayes estimation of reliability for the two-parameter lognormal distribution. Communications in Statistics: Theory and Methods, 6(5):443–457, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Passi:1978:SSR

[PW78] Ranjit M. Passi and James S. Williams. Sample-size requirements for confidence bounds of prescribed precision on multivariate normal quantiles. Communications in Statistics: Theory and Methods, 7(8): 739–753, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Padgett:1979:ETP

[PW79] W. J. Padgett and L. J. Wei. Estimation for the three-parameter inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 8(2):129–137, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pathak:1979:NPE

[PZW79] Pramod K. Pathak, William J. Zimmer, and Raymond E. Williams. Non-parametric estimation of an acceleration parameter. Communications in Statistics: Theory and Methods, 8(4):367–383, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quasenberry:1976:OTS

[QS76] C. P. Quasenberry and R. R. Starbuck. On optimal tests for separate hypotheses and conditional probability integral transformations. Communications in Statistics: Theory and Methods, 5(6): 507–524, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quesenberry:1978:SIP

[Que78] C. P. Quesenberry. On similarity and independence properties of composite EDF goodness-of-fit tests. Communications in Statistics: Theory and Methods, 7(8):717–723, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramberg:1977:SBP

[Ram77a] John S. Ramberg. Selecting the best predictor variate. Communications in Statistics: Theory and Methods, 6(11):1133–1147, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramig:1977:GWB

[Ram77b] P. F. Ramig. On the Galton–Watson branching process: application of a method of moments and recursive determination for certain distributions. Communications in Statistics: Theory and Methods, 6(2):175–196, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1978:CBL

[Rao78a] C. Radhakrishna Rao. Choice of best linear in the Gauss–Markoff model with a singular dispersion matrix. Communications in Statistics: Theory and Methods, 7(13):1199–1208, 1978. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1978:NUT

[Rao78b] C. Radhakrishna Rao. A note on the unified theory of least squares. Communications in Statistics: Theory and Methods, 7(5):409–411, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1979:EPS

[Rao79] C. Radhakrishna Rao. Estimation of parameters in the singular Gauss–Markoff model. Communications in Statistics: Theory and Methods, 8(14):1353–1358, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ray:1976:NTT

[Ray76] R. M. Ray. A new C(α) test for 2 × 2 tables. Communications in Statistics: Theory and Methods, 5(6):545–563, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1978:TMP

[RC78] Poduri S. R. S. Rao and Yogendra P. Chaubey. Three modifications of the principle of the MINQUE. Communications in Statistics: Theory and Methods, 7(8):767–778, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Read:1977:PCS

[Rea77] Campbell B. Read. Partitioning chi-square in contingency tables: a teaching approach. Communications in Statistics: Theory and Methods, 6(6):553–562, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Resek:1976:EPG

[Res76] Robert W. Resek. Estimation of the parameters of a general Student’s t distribution. Communications in Statistics: Theory and Methods, 5(7):635–645, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reed:1979:STS

[RF79] Allen H. Reed and Mary E. Frantz. A sequential two sample t test using Welch type modification for unequal variances. Communications in Statistics: Theory and Methods, 8(14):1459–1471, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rinott:1978:TSS

[Rin78] Yosef Rinott. On two-stage selection procedures and related probability-inequalities. Communications in Statistics: Theory and Methods, 7(8):799–811, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raghavarao:1979:ACA

[RN79] D. Raghavarao and G. Nageswararao. Analysis of covariance as an alternative approach to the analysis of non-orthogonal lay-outs. Communications in Statistics: Theory and Methods, 8(8):843–854, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roberts:1976:WUT

[Rob76] Harry V. Roberts. For what use are tests of hypotheses and tests of significance. Communications in Statistics: Theory and Methods, 5(8):753–761, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robinson:1977:FPA

[Rob77] P. M. Robinson. Finite-parameter approximations to frequency response function. Communications in Statistics: Theory and Methods, 6(7):661–677, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rolph:1976:CSE

[Rol76] John E. Rolph. Choosing shrinkage estimators for regression problems. Communications in Statistics: Theory and Methods, 5(9): 789–802, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1976:ERP

[RR76] P. V. Rao and C. Littell Ramon. An estimator of relative potency. Communications in Statistics: Theory and Methods, 5(2):183–189, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruben:1976:ESD

[Rub76] Harold Ruben. On the evaluation of Steck’s determinant for rectangle probabilities of uniform order statistics. Communications in Statistics: Theory and Methods, 5(6):535–543, 1976. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Russell:1976:COC

[Rus76] K. G. Russell. The connectedness and optimality of a class of rowcolumn designs. Communications in Statistics: Theory and Methods, 5(15):1479–1488, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reed:1977:EBC

[RW77] Allen H. Reed and Gaw T. Wu. Estimation of bias in classical linear regression slope when the proper model is functional linear regression. Communications in Statistics: Theory and Methods, 6 (5):405–416, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rogers:1977:EML

[RY77] G. S. Rogers and D. L. Young. Explicit maximum likelihood estimators for certain patterned covariance matrices. Communications in Statistics: Theory and Methods, 6(2):121–133, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Santner:1976:TSP

[San76] Thomas J. Santner. A two stage procedure for selecting δ^{∗}-optimal means in the normal model. Communications in Statistics: Theory and Methods, 5(3):283–292, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saw:1977:ICR

[Saw77] John G. Saw. On inequalities in constrained random variables. Communications in Statistics: Theory and Methods, 6(13):1301–1304, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schucany:1976:AMS

[SB76a] William R. Schucany and James Beckett. Analysis of multiple sets of incomplete rankings. Communications in Statistics: Theory and Methods, 5(14):1327–1334, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1976:CTG

[SB76b] Robert M. Smith and Lee J. Bain. Correlation type goodness-of-fit statistics with censored sampling. Communications in Statistics: Theory and Methods, 5(2):119–132, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scheaffer:1976:NAC

[Sch76] R. L. Scheaffer. A note on approximate confidence limits for the negative binomial model. Communications in Statistics: Theory and Methods, 5(2):149–158, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schey:1977:ADO

[Sch77] H. M. Schey. The asymptotic distribution of the one-sided Kolmogorov–Smirnov statistic for truncated data. Communications in Statistics: Theory and Methods, 6(14):1361–1366, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sclove:1977:PMM

[Scl77] Stanley L. Sclove. Population mixture models and clustering algorithms. Communications in Statistics: Theory and Methods, 6(5): 417–434, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:1976:MSR

[Sco76] Stuart Scott. Moments of suprema of random variables. Communications in Statistics: Theory and Methods, 5(7):673–680, 1976. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shuster:1976:QRA

[SD76] J. J. Shuster and F. H. Dietrich. Quantal response assays by inverse regression. Communications in Statistics: Theory and Methods, 5 (4):293–305, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1977:FGS

[SD77] A. K. Md. Ehsanes Saleh and J.-P. Dionne. On a further generalization of the savage test. Communications in Statistics: Theory and Methods, 6(12):1213–1221, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Searle:1979:ACM

[Sea79] S. R. Searle. Alternative covariance models for the 2-way crossed classification. Communications in Statistics: Theory and Methods, 8(8):799–818, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seely:1978:CSS

[See78] Justus Seely. A complete sufficient statistic for the linear model under normality and a singular covariance matrix. Communications in Statistics: Theory and Methods, 7(15):1465–1473, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seheult:1978:MBL

[Seh78] Allan Seheult. Minium bias or least squares estimation? Communications in Statistics: Theory and Methods, 7(3):277–283, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1978:EPM

[Sen78a] A. R. Sen. Estimation of the population mean when the coefficient of variation is known. Communications in Statistics: Theory and Methods, 7(7):657–672, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1978:TSS

[Sen78b] Pranab Kumar Sen. Time-sequential statistical procedures: a preface. Communications in Statistics: Theory and Methods, 7(4):311– 314, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1979:NRS

[Sen79] Pranab Kumar Sen. Nonparametric repeated significance tests for some analysis of covariance models. Communications in Statistics: Theory and Methods, 8(8):819–841, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinlia:1976:BIA

[SG76] S. K. Sinlia and Irwin Guttman. Bayesian inference about the reliability function for the exponential distributions. Communications in Statistics: Theory and Methods, 5(5):471–479, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1977:BFD

[SG77] J. N. Srivastava and S. Ghosh. Balanced 2^{m} factorial designs of resolution V which allow search and estimation of one extra unknown effect, 4 ≤ m ≤ 8. Communications in Statistics: Theory and Methods, 6(2):141–166, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1978:MLA

[SH78] David W. Smith and R. R. Hocking. Maximum likelihood analysis of the mixed model: the balanced case. Communications in Statistics: Theory and Methods, 7(13):1253–1266, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaked:1977:SIC

[Sha77] Moshe Shaked. Statistical inference for a class of life distributions. Communications in Statistics: Theory and Methods, 6(13):1323– 1339, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaked:1979:EGH

[Sha79] Moshe Shaked. An estimator for the generalized hazard rate function. Communications in Statistics: Theory and Methods, 8(1):17– 33, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shinozaki:1978:NEC

[Shi78] N. Shinozaki. A note on estimating the common mean of k normal distributions and the Stein problem. Communications in Statistics: Theory and Methods, 7(15):1421–1432, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shuster:1976:PED

[Shu76] J. J. Shuster. The precision of an estimator due to Ryan. Communications in Statistics: Theory and Methods, 5(9):867–874, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Simon:1977:UCI

[SHW77] Richard Simon, David G. Hoel, and George H. Weiss. The use of covariate information in the sequential analysis of dichotomous response experiments. Communications in Statistics: Theory and Methods, 6(8):777–788, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sia:1977:SPC

[Sia77] L. L. Sia. Some properties of connectedness in two-way designs. Communications in Statistics: Theory and Methods, 6(12):1165– 1169, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schrader:1977:RAV

[SK77] Ronald M. Schrader and Joseph W. Mc Kean. Robust analysis of variance. Communications in Statistics: Theory and Methods, 6(9): 879–894, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Skillings:1978:ACI

[Ski78] John H. Skillings. Adaptively combining independent Jonckheere statistics in a randomized block design with unequal scales. Communications in Statistics: Theory and Methods, 7(11):1027–1039, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salama:1978:EMP

[SKT78] Ibrahim A. Salama, Gary G. Koch, and Dennis H. Tolley. On the estimation of the most probable number in a serial dilution experiment. Communications in Statistics: Theory and Methods, 7(13): 1267–1281, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1978:FWS

[SL78] J. W. H. Swanepoel and F. Lombard. Fixed width sequential confidence interval for the median of a distribution. Communications in Statistics: Theory and Methods, 7(9):829–835, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1979:BPT

[SL79] Y. S. Sathe and S. R. Lingras. Bounds for the Pearson type IV tail probabilities. Communications in Statistics: Theory and Methods, 8(6):533–541, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1976:MAR

[SM76] P. A. V. B. Swamy and J. S. Mehta. Minimum average risk estimators for coefficients in linear models. Communications in Statistics: Theory and Methods, 5(9):803–818, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1977:NMA

[SM77] P. A. V. B. Swamy and J. S. Mehta. A note on minimum average risk estimators for coefficients in linear models. Communications in Statistics: Theory and Methods, 6(12):1181–1186, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1977:NTB

[Smi77a] Paul J. Smith. A nonparametric test for bivariate circular symmetry based on the empirical cdf. Communications in Statistics: Theory and Methods, 6(3):209–220, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1977:SRI

[Smi77b] Robert M. Smith. Some results on interval estimation for the twoparameter Weibull or extreme-value distribution. Communications in Statistics: Theory and Methods, 6(13):1311–1321, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1979:EVA

[Smi79] E. J. Smith. An experimenter’s view of the application of statistics to cloud-seeding experiments. Communications in Statistics: Theory and Methods, 8(10):223–230, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1978:FR

[SMR78a] P. A. V. B. Swamy, J. S. Mehta, and P. N. Rappoport. Further reply. Communications in Statistics: Theory and Methods, 7(12): 1171–1172, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1978:R

[SMR78b] P. A. V. B. Swamy, J. S. Mehta, and P. N. Rappoport. Reply. Communications in Statistics: Theory and Methods, 7(12):1163– 1168, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swamy:1978:TME

[SMR78c] P. A. V. B. Swamy, J. S. Mehta, and P. N. Rappoport. Two methods of evaluating Hoerl and Kennard’s ridge regression. Communications in Statistics: Theory and Methods, 7(12):1133–1155, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Snee:1979:EDM

[Sne79] Ronald D. Snee. Experimental designs for mixture systems with multicomponent constraints. Communications in Statistics: Theory and Methods, 8(4):303–326, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Susarla:1979:EBI

[SO79] V. Susarla and T. O’Bryan. Empirical Bayes interval estimates involving uniform distributions. Communications in Statistics: Theory and Methods, 8(4):385–397, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shah:1976:AMP

[SP76a] K. R. Shah and S. C. Puri. Application of MINQUE procedure to block designs. Communications in Statistics: Theory and Methods, 5(2):191–196, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Susarlam:1976:EBT

[SP76b] V. Susarlam and E. G. Phadia. Empirical Bayes testing of a distribution function with Dirichlet process priors. Communications in Statistics: Theory and Methods, 5(5):455–469, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1976:DST

[Spu76] John D. Spurrier. Double sample tests for the equality of two means. Communications in Statistics: Theory and Methods, 5(4):343–349, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1978:TSS

[Spu78] John D. Spurrier. Two stage sign tests that allow first stage acceptance and rejection. Communications in Statistics: Theory and Methods, 7(4):399–408, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1979:EME

[SR79] V. K. Srivastava and Baldev Raj. The existence of the mean of the estimator in seemingly unrelated regressions. Communications in Statistics: Theory and Methods, 8(7):713–717, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1978:BFP

[SRC78] Bimal Kumar Sinha, B. Raja Rao, and Bernard Clement. Behrens– Fisher problem under the assumption of homogeneous coefficients of variation. Communications in Statistics: Theory and Methods, 7(7):637–656, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1976:CDM

[SS76] Bimal Kumar Sinha and Bikas Kumar Sinha. On a characterization of the dispersion matrix based on the properties of regression.

Communications in Statistics: Theory and Methods, 5(13):1215– 1224, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1979:ISS

[SS79a] Y. S. Sathe and R. G. Shenoy. On inequalities for statistics similar to Student’s t. Communications in Statistics: Theory and Methods, 8(6):575–579, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1979:PCT

[SS79b] Agam Nath Sinha and Pranab Kumar Sen. Progressively censored tests for clinical experiments and life testing problems based on weighted empirical distributions. Communications in Statistics: Theory and Methods, 8(9):871–897, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiyama:1976:SUD

[ST76] T. Sugiyama and H. Tong. On a statistic useful in dimensionality reduction in multivariable linear stochastic system. Communications in Statistics: Theory and Methods, 5(8):711–721, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steece:1979:BAM

[Ste79] Bert M. Steece. A Bayesian approach to model adequacy. Communications in Statistics: Theory and Methods, 8(14):1393–1402, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stokes:1977:RSS

[Sto77] S. Lynne Stokes. Ranked set sampling with concomitant variables. Communications in Statistics: Theory and Methods, 6(12):1207– 1211, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvey:1978:AOD

[STT78] S. D. Silvey, D. H. Titterington, and B. Torsney. An algorithm for optimal designs on a design space. Communications in Statistics: Theory and Methods, 7(14):1379–1389, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Studden:1978:DLD

[Stu78] W. J. Studden. Designs for large degree polynomial regression. Communications in Statistics: Theory and Methods, 7(14):1391–1397, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Subrahmaniam:1978:PPD

[Sub78] Kathleen Subrahmaniam. The Pascal–Poisson distribution revisited: estimation and efficiency. Communications in Statistics: Theory and Methods, 7(7):673–683, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1978:FAD

[Sug78] Nariaki Sugiura. Further analysts of the data by Akaike’s information criterion and the finite corrections. Communications in Statistics: Theory and Methods, 7(1):13–26, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schucany:1977:ADF

[SW77a] William R. Schucany and Wayne A. Woodward. Adjusting the degrees of freedom for the jackknife. Communications in Statistics: Theory and Methods, 6(5):439–442, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Skillings:1977:TOA

[SW77b] John H. Skillings and Douglas A. Wolfe. Testing for ordered alternatives by combining independent distribution-free block statistics. Communications in Statistics: Theory and Methods, 6(15):1453– 1463, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swindel:1976:GRE

[Swi76] Benee F. Swindel. Good ridge estimators based on prior information. Communications in Statistics: Theory and Methods, 5(11):1065– 1075, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tamhane:1977:CAM

[Tam77a] Ajit C. Tamhane. Corrections and amendments multiple comparisons in Model I one-way ANOVA with unequal variances. Communications in Statistics: Theory and Methods, 6(15):1535–1536, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tamhane:1977:MCM

[Tam77b] Ajit C. Tamhane. Multiple comparisons in Model I one-way ANOVA with unequal variances. Communications in Statistics: Theory and Methods, 6(1):15–32, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:1976:SMT

[TB76] Peter Tan and B. Bernholtz. On structural models and a theorem of D. Basu. Communications in Statistics: Theory and Methods, 5 (13):1245–1256, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tamhane:1977:TSM

[TB77] Ajit C. Tamhane and Robert E. Bechhofer. A two-stage minimax procedure with screening for selecting the largest normal mean. Communications in Statistics: Theory and Methods, 6(11):1003– 1033, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tamhane:1979:TSM

[TB79] Ajit C. Tamhane and Robert E. Bechhofer. A two-stage minimax procedure with screening for selecting the largest normal mean (II): an improved PCS lower bound and associated tables. Communications in Statistics: Theory and Methods, 8(4):337–358, 1979. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tubbs:1976:ESS

[TC76] J. D. Tubbs and W. A. Coberly. An empirical sensitivity study of mixture proportion estimators. Communications in Statistics: Theory and Methods, 5(12):1115–1125, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Talwar:1977:RTH

[TG77] Prem P. Talwar and James E. Gentle. A robust test for the homogeneity of scales. Communications in Statistics: Theory and Methods, 6(4):363–369, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taylor:1979:UTA

[TG79a] M. S. Taylor and F. E. Grubbs. Use of tail areas for testing goodnessof-fit. Communications in Statistics: Theory and Methods, 8(15): 1575–1584, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1979:SAK

[TG79b] Ram C. Tripathi and John Gurland. Some aspects of the Kemp families of distributions. Communications in Statistics: Theory and Methods, 8(9):855–869, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1975:NST

[Tik75] M. L. Tiku. A new statistic for testing suspected outliers. Communications in Statistics, 4(8):737–752, 1975. CODEN COSTCX. ISSN 0090-3272 (print), 2332-3965 (electronic). See comment [Haw77] and rejoinder [Tik77].

Tiku:1977:RCN

[Tik77] M. L. Tiku. Rejoinder: “Comment on ‘A new Statistic for Testing Suspected Outliers’ ”. Communications in Statistics: Theory and Methods, 6(14):1417–1422, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:1977:API

[Ton77] Y. L. Tong. Applications of a probability inequality to ranking and selection and other related problems. Communications in Statistics: Theory and Methods, 6(11):1105–1120, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:1979:CRB

[Ton79] Y. L. Tong. Counterexamples to a result of Broemeling on simultaneous inferences for variance ratios of some mixed linear models. Communications in Statistics: Theory and Methods, 8(12):1197– 1204, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1977:ETG

[TS77] M. L. Tiku and D. E. Stewart. Estimating and testing group effects from Type I censored normal samples in experimental design. Communications in Statistics: Theory and Methods, 6(15):1485–1501, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Turnbull:1976:MDR

[Tur76] Bruce W. Turnbull. Multiple decision rules for comparing several populations with a fixed known standard. Communications in Statistics: Theory and Methods, 5(13):1225–1244, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ury:1979:DFC

[UW79] Hans K. Ury and Alvin D. Wiggins. Distribution-free confidence bounds for Pr(y < x) when F (x) and G(y) = F (x−c) are continuous and symmetric. Communications in Statistics: Theory and Methods, 8(12):1247–1253, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinod:1978:ERE

[Vin78a] H. D. Vinod. Equivariance of ridge estimators through standardization, a note. Communications in Statistics: Theory and Methods, 7 (12):1157–1161, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinod:1978:R

[Vin78b] H. D. Vinod. A rejoinder. Communications in Statistics: Theory and Methods, 7(12):1169–1170, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VanRyzin:1978:MSE

[VW78] J. Van Ryzin and H. C. Wang. The methods for smooth estimation of discrete distributions. Communications in Statistics: Theory and Methods, 7(3):211–228, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Watkins:1979:EFM

[Wat79] T. A. Watkins. On estimating a function of the mean of a multivariate normal distribution. Communications in Statistics: Theory and Methods, 8(3):245–256, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williford:1979:BEP

[WB79] W. O. Williford and S. K. Bingham. Bayesian estimation of the parameters in two modified Poisson distributions. Communications in Statistics: Theory and Methods, 8(13):1315–1326, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wegman:1977:MCS

[WC77] Edward J. Wegman and Raymond J. Carroll. A Monte Carlo study of robust estimators of location. Communications in Statistics: Theory and Methods, 6(9):795–812, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wackerly:1976:PCM

[WD76] D. D. Wackerly and F. H. Dietrich. Pairwise comparisons of matched proportions. Communications in Statistics: Theory and Methods, 5 (15):1455–1467, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weiss:1976:TST

[Wei76] Lionel Weiss. Two-sample tests and tests of fit. Communications in Statistics: Theory and Methods, 5(13):1275–1285, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:1978:CTA

[Wei78a] L. J. Wei. A class of treatment assignment rules for sequential experiments. Communications in Statistics: Theory and Methods, 7(3):285–295, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weiss:1978:STW

[Wei78b] Lionel Weiss. Some tests of whether several samples are from identical populations. Communications in Statistics: Theory and Methods, 7(12):1099–1127, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williams:1976:TAB

[Wil76] George W. Williams. Testing the association between health status and an environmental index. Communications in Statistics: Theory and Methods, 5(5):421–427, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williams:1978:SDD

[Wil78] E. J. Williams. A simple derivation of the distribution of the multiple correlation coefficient. Communications in Statistics: Theory and Methods, 7(15):1413–1420, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woolson:1976:MCI

[WLCC76] R. F. Woolson, J. D. Leeper, J. W. L. Cole, and W. R. Clarke. A Monte Carlo investigation of a statistic for a bivariate missing data problem. Communications in Statistics: Theory and Methods, 5(7):681–688, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wiorkowski:1979:P

[WO79] John J. Wiorkowski and Patrick L. Odell. Preface. Communications in Statistics: Theory and Methods, 8(10):953–954, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wood:1978:NHL

[Woo78] Constance L. Wood. On null-hypothesis limiting distributions of Kolmogorov–Smiirnov type statistics with estimated location and scale parameters. Communications in Statistics: Theory and Methods, 7(12):1181–1198, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weiner:1977:SMA

[WP77] Dan Weiner and Peter Purdue. A semi-Markov approach to stochastic compartmental models. Communications in Statistics: Theory and Methods, 6(12):1231–1243, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1978:SIP

[Wu78] Chien-Fu Wu. Some iterative procedures for generating nonsingular optimal designs. Communications in Statistics: Theory and Methods, 7(14):1399–1412, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wynn:1977:CCC

[Wyn77] H. P. Wynn. The combinatorial characterization of certain connected 2 × j × k three-way layouts. Communications in Statistics: Theory and Methods, 6(10):945–953, 1977. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1977:RBL

[Yam77] Hajime Yamato. Relations between limiting Bayes estimates and the U-statistics for estimable parameters of degrees 2 and 3. Communications in Statistics: Theory and Methods, 6(1):55–66, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1978:UPI

[YBW78] P. M. V. Yu, T. L. Bouillon, and I. A. Watkins. The use of prior information in linear regression. Communications in Statistics: Theory and Methods, 7(1):81–95, 1978. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yohai:1976:LEB

[YM76] Victor J. Yohai and Ricardo A. Maronna. Location estimators based on linear combinations of modified order statistics. Communications in Statistics: Theory and Methods, 5(5):481–486, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuen:1977:RSS

[Yue77] Karen K. Yuen. Robustness of some sequential procedures. Communications in Statistics: Theory and Methods, 6(1):43–54, 1977. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zerbe:1979:RAR

[Zer79] O. Zerbe. Randomization analysis of randomized blocks design extended to growth and response curves. Communications in Statistics: Theory and Methods, 8(2):191–205, 1979. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zacks:1976:TEI

[ZS76] S. Zacks and H. Solomon. On testing and estimating the interaction between treatments and environmental conditions in binomial experiments: The case of two stations. Communications in Statistics: Theory and Methods, 5(3):197–223, 1976. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).